### v1

# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

Abstract.
Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation. To address such concerns, numerous AI-generated Image (AIGI) Detectors have been proposed and achieved promising performance in identifying fake images. However, there still lacks a systematic understanding of the adversarial robustness of AIGI detectors. In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far. For the task of AIGI detection, we propose a new attack containing two main parts. First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution. Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models, e.g. transferring adversarial examples across CNNs and ViTs. This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training. We name our method as frequency-based post-train Bayesian attack, or FPBA. Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario. The code will be shared upon acceptance.
AI-generated Image Detection, Adversarial Examples
ccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Computer vision
1. Introduction
The notable progress in generative models, such as GANs (Goodfellow et al., 2014a) and Diffusion models (Ho et al., 2020), is driving the flourishing development of the image synthesis domain. These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images. Moreover, a variety of free and open-source tools facilitates the effortless creation of fake images. However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
Consequently, numerous detectors have been proposed(Durall et al., 2020; Chai et al., 2020; Corvi et al., 2023b) to identify AI-generated images (AIGI). Recent state-of-the-art detectors (Wang et al., 2020; Zhong et al., 2023) rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models. However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real. Despite several works (Carlini and Farid, 2020; Hussain et al., 2021; Neekhara et al., 2021; Hou et al., 2023; Jia et al., 2022a) exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored. Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs. These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
Figure 1. Illustration of the proposed method.
In this paper, we show that AIGI detectors are vulnerable to adversarial attacks. Considering that many works (Dzanic et al., 2020; Frank et al., 2020) have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain. As illustrated in Fig. 3, the frequency components that different detectors focus on significantly vary from each other. Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models. To further improve the adversarial transferability across heterogeneous models, e.g. transferring adversarial examples across Visual Transformers (ViTs) and Convolutional Neural Networks (CNNs), we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate. In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process. As a result, we propose the first adversarial attack for general AIGI detection, to add adversarial perturbations in the frequency domain from a post-train Bayesian perspective. We name our method Frequency-based Post-train Bayesian Attack, or FPBA. A high-level illustration of our method and the key differences between our method and the previous method are shown in Fig. 1.
The contributions of this work can be summarized as follows:
  • To our best knowledge, we are the first to systemically investigate the adversarial vulnerability of state-of-the-art AIGI detectors. We extensively evaluate the adversarial robustness of traditionally trained models, defense models, and cross-generated detection within real-world scenarios. We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
  • We propose a novel attack against AIGI detection by exploring the vulnerable frequency region in a Bayesian manner. A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
  • We conduct comprehensive experiments on eight spatial-based detectors with CNN-based, Vit-based and CLIP-based architectures, and two frequency-based detectors using AIGI and Deepfake datasets. The extensive results demonstrate our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
2. Related Work
AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs (Goodfellow et al., 2014a), VAEs (Kingma and Welling, 2014), and Diffusion models (Ho et al., 2020), has brought about substantial risks in the dissemination of disinformation. Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods to primarily focus on face forgeries. However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects. Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances. Data-driven-based approaches (Durall et al., 2020; Chai et al., 2020; Wang et al., 2020; Corvi et al., 2023b) have demonstrated promising results on generated data identically distributed. However, their performance significantly degraded when presented with previously unseen generated images. More recently, certain methods (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) focus on analyzing forgery traces in the frequency domain and learning relevant forgery features (Corvi et al., 2023a; Liu et al., 2023; Pontorno et al., 2024; Zhang et al., 2019) to enhance the generalization capability of the detection system. In pursuit of these objectives, some methods have been proposed to learn local forgery features (Ju et al., 2023; Lorenz et al., 2023) in the spatial domain. AIGI fingerprint representations (Jeong et al., 2022) are acquired through the learning of noise patterns. Notably, studies such as (Liu et al., 2022; Zhong et al., 2023) have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance. A sequence of detection methodologies (Tan et al., 2023; Liu et al., 2023; Cozzolino et al., 2023; Chang et al., 2023) has been progressively investigated, relying on the practice of freezing pre-trained models. These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers. For instance, UniFD (Ojha et al., 2023) explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection. Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error (Wang et al., 2023a; Ricker et al., 2024), adversarial teacher-student discrepancy-aware framework (Zhu et al., 2023a) or diffusion noise(Zhang and Xu, 2023).
The existing models consistently demonstrate enhancement in detecting AI-generated images. However, their behaviors under adversarial attacks have not been thoroughly examined. Our work is complementary to existing research by making the first attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly. The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including face recognition (Jia et al., 2022b), autonomous driving (Zhu et al., 2023b; Xie et al., 2023) and human activity recognition (Diao et al., 2021). Very recently, the adversarial attack against Deepfake detection has drawn attention. Hussain et al. (Hussain et al., 2021) employ gradient sign-based methods (Kurakin et al., 2018; Goodfellow et al., 2014b) to evaluate the adversarial robustness of video Deepfake detectors. Li et al. (Li et al., 2021) generate adversarial fake images along the latent face manifold. Carlini et al. (Carlini and Farid, 2020) and Neekhara et al. (Neekhara et al., 2021) investigate the adversarial threat to Deepfake detection under the black-box setting. Jia et al. (Jia et al., 2022a) propose to generate adversarial perturbations in the frequency domain. Hou et al. (Hou et al., 2023) narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation. However, the adversarial robustness of AIGI detection has received little attention. In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
3. Methodology
Figure 2. The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
3.1. Preliminaries
Let $\mathbf{x}$ and $y$ represent the original image and its corresponding label. $f_{\Theta}$ denotes the AI-generated image detectors. We aim to inject adversarial perturbation into the original image that makes the detector misclassify. Such an adversary problem can be optimized by minimizing the predictive probability, i.e. maximizing the classification loss:
(1) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
where $L$ is the binary cross entropy loss in AI-generated image detection. Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget. Eq. 1 can be performed with iterative gradient-based methods, such as PGD (Madry et al., 2017) or I-FGSM (Kurakin et al., 2018):
(2) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
(3) $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$
3.2. Frequency-based Analysis and Attacks
[…]
Many AIGI detection approaches distinguish between real and fake images via subtle artifacts (Wang et al., 2020; Jeong et al., 2022; Zhong et al., 2023). While these subtle clues are invisible in the spatial domain, a series of works (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) demonstrate that there are obvious differences between real and fake images in the frequency domain. This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective. To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain. To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map (Long et al., 2022) to visualize the sensitive components of real and fake images across different models:
(4) $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})}$
[…]
(5) $\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$
where $\Gamma(\cdot)$ denotes the random spectrum transformation (Long et al., 2022). $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$. As shown in Fig. 3(a), tuning the spectrum saliency map can cover most of the other models. We hence conduct adversarial attack in the frequency domain via spectrum transformation:
(6) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
3.3. Exploring the Surrogate Posterior Space
Although tuning the spectrum transformation in Eq. 6 can simulate different substitute models with a homogeneous architecture (Long et al., 2022), it shows limited transferability when applied to heterogeneous architectures, e.g. transferring adversarial examples across ViTs and CNNs. This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e. exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models. Therefore, we redefine Eq. 6 by minimizing the Bayesian posterior predictive distribution:
$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,p(y\mid\Gamma(\tilde{\mathbf{x}}),\mathcal{D})$
(7) $=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
where $p(\Theta\mid\mathcal{D})\propto p(\mathcal{D}\mid\Theta)p(\Theta)$. $\mathcal{D}$ is the dataset and $p(\Theta)$ is the prior of model weights. Attacking Bayesian Neural Networks (BNNs) rather than a single DNN allows for the output fusion from an ensemble of infinitely many DNNs with diverse predictions, thereby improving adversarial transferability.
Figure 4. The architecture of the appended model. $\sigma$ means the sigmoid layer.
3.3.1. Post-train Bayesian Strategy for Frequency-based Attack
However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors. First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs (Izmailov et al., 2021). Hence computing and sampling the posterior distribution is an intractable problem. Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space. Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets (Zhu et al., 2024; He et al., 2021). From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
Therefore, we propose a post-train Bayesian strategy to turn a single surrogate into a Bayesian one, without the need for re-training. The parameters over the pre-trained surrogate are represented as $\Theta=[\theta,\theta_{c}]$, in which $f_{\theta}$ represents the feature extraction backbone, and $f_{\theta_{c}}$ represents the fully-connected layer network for classification. As shown in Fig. 4, we fix the pre-trained surrogate and append a tiny Bayesian component $g_{\theta^{\prime}}$ behind the feature extraction backbone $f_{\theta}$. The new logits can be computed via a skip connection:
(8) $\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x})$
We choose to apply Bayesian Model Averaging to optimize the appended Bayesian model:
(9) $\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)$
where $K$ is the number of appended models. $\Theta$ is fixed to avoid re-training. We surprisingly find that adopting a simple MLP layer for appended models works well in all cases, hence training the appended models is much faster than re-training a surrogate. Finally, the frequency-based post-train Bayesian attack can be conducted with iterative gradient-based methods:
(10) $\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}$
3.3.2. Inference on Bayesian Appended Models
$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo(Springenberg et al., 2016) to sample appended model $\theta^{\prime}$ in each iteration:
$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I})$
(11) $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2}$
where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically (Springenberg et al., 2016).
3.4. Hybrid Adversarial Attack
Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain (Wang et al., 2020; Jeong et al., 2022). We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains. Specifically, we define the hybrid attack as:
(12) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\}$
(13) $g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i}$
(14) $d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k})$
where $g^{i}_{k}$ and $d^{i}_{k}$ are the gradients computed in the frequency domain and spatial domain respectively. For frequency gradient, we conduct random spectrum transformation with $N$ times to get more diverse spectrums. Our proposed method leverages both spatial attack gradients and frequency attack gradients in a Bayesian manner, aiming to further narrow the discrepancy between surrogate models and victim models. The complete algorithm of our method is presented in Algorithm 1. An overview illustration of FPBA is shown in Fig. 2.
Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models; ;
Output: The adversarial example $\tilde{\mathbf{x}}$;
// Post-train Bayesian Optimization
Randomly initialize $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
for j = 1 to $N_{tra}$ do
       for n = 1 to $K$ do
             Randomly sample a mini-batch data $\{\mathbf{x},y\}_{j}$;
             Compute $\mathbf{h}_{\theta^{\prime}_{k}}=\frac{\partial logp(y|\mathbf{x},\Theta,\theta^{\prime}_{k})}{\partial\theta^{\prime}_{k}}$;
             for t = 1 to $M_{\theta^{\prime}}$ do
                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Sec. 3.3.2;
             end for
       end for
end for
return $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$;
// Frequency-based Post-train Bayesian Attack
$\tilde{\mathbf{x}}^{0}=\mathbf{x}$ ;
for i = 1 to $I$ do
       for n = 1 to $N$ do
             Get spectrum transformation output $\Gamma(\tilde{\mathbf{x}}^{i})$ using Eq. 5 ;
       end for
      for k = 1 to $K$ do
             Averaging frequency gradient $g_{k}$ using Eq. 13 ;
             Calculating spatial gradient $d_{k}$ using Eq. 14 ;
       end for
      Sample $\tilde{\mathbf{x}}^{i+1}$ from $\tilde{\mathbf{x}}^{i}$ via Eq. 12 ;
end for
return $\tilde{\mathbf{x}}$;
Algorithm 1 Inference on FPBA
4. Experimets
4.1. Experimental Settings
Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection. Synthetic LSUN is a commonly used dataset proposed by CNNSpot (Wang et al., 2020), containing 720k real images from LSUN and 360k fake images generated by ProGAN (Karras et al., 2018). GenImage (Zhu et al., 2024) is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models. Following the protocol in (Zhu et al., 2024), we employ a subset of GenImage collecting 162k real images from Imagenet (Deng et al., 2009) and 162k Stable Diffusion(SD) V1.4 (Rombach et al., 2022) generated images for training. The images generated by the other generators are used for testing in Sec. 4.3.2. To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by (Shamshad et al., 2023). The Deepfake dataset consists of 50k real face images from FFHQ (Karras et al., 2019) and 50k generated face images generated by StyleGAN2 (Karras et al., 2019). After training, we collect only the correctly classified testing samples for attack in evaluation.
Evaluated Models: We extensively evaluate the transferability of adversarial examples by eight different models across heterogeneous architectures under the spatial and frequency domain. We choose spatial-based classification models ResNet-50 (He et al., 2016) and MobileNet (Howard et al., 2017) as the source models. For target models, apart from the above models, we also use spatial-based models including EfficientNet (Tan and Le, 2019), Vision Transformer(Vit) (Dosovitskiy et al., 2020) and Swin-Transformer(Swin-Vit) (Liu et al., 2021), and frequency-based detectors DCTA (Frank et al., 2020) and Spec (Zhang et al., 2019). We report the accuracy of these detectors in the supplement. Besides, we also evaluate the robustness of state-of-the-art AIGI detection methods, including GramNet (Liu et al., 2020), LGrad (Tan et al., 2023), LNP (Liu et al., 2022) and UnivFD (Ojha et al., 2023).
Compared Methods: We adopt gradient-based methods, I-FGSM (Kurakin et al., 2018), PGD (Madry et al., 2017), MI (Dong et al., 2018) and Frequency-based methods S2(Long et al., 2022). We also consider ensemble-based methods Ensemble (Dong et al., 2018) and SVRE (Xiong et al., 2022). For a fair comparison, we ran 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all attacks.
Implementation Details: We choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001. We follow the same data augmentation strategy in CNNSpot (Wang et al., 2020). Before cropping, images are blurred with $\sigma$ Uniform[0, 3] with 10% probability, and JPEG-ed with 10% probability. We transformed images to 224 pixels and apply ImageNet normalization. For post-train Bayesian optimization, we follow the default setting in (Wang et al., 2023b). Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate. Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$. For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following (Long et al., 2022). All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
Table 1. The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN, GenImage subset and Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
ResNet PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
MobileNet PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
IFGSM 78.3 71.5 30.9 34.0 27.1 28.3 11.0 27.8 32.9
MIFGSM 78.3 75.3 30.1 35.8 27.8 28.3 15.0 28.2 34.4
ResNet-50 PGD 99.3 88.0 48.5 52.2 48.6 49.9 15.4 48.1 50.1
S2I 98.0 91.0 66.8 64.7 36.7 50.1 16.9 53.0 54.2
FPBA(Ours) 100 96.6 59.4 67.3 47.9 49.8 21.1 51.2 56.2
IFGSM 9.0 8.9 8.9 58.9 9.1 9.1 4.4 8.5 8.3
MIFGSM 9.7 9.2 8.9 59.0 9.5 9.0 5.9 8.8 8.7
MobileNet PGD 42.3 39.9 36.4 97.1 48.4 49.8 14.2 40.0 38.7
S2I 35.6 33.1 26.8 78.7 24.1 43.4 6.2 28.0 28.2
GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
ResNet PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
MobileNet PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
4.2. The Evaluation on Normally Trained Models
We show the attack performance against all models on all datasets in Tab. 1. Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods. Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.8%, while S2I, PGD, IFGSM and MIFGSM only have 95.6%, 95.4%, 65.4% and 65.5%. Under the black-box setting, FPBA still achieves the highest average transfer success rate of 55.3%, surpassing MIFGSM, PGD and S2I by a margin of 27.7%, 7.6% and 8.4% respectively.
Next, We find that the attack results for FPBA are usually not the best on frequency-based models. To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq. 4 in Fig. 5. Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig. 3), the frequency components focused on in frequency-based models are very sparse. Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain. Nevertheless, our method still has the best average success rate over all models. In addition, we combine spatial-based gradient and frequency-based to further improve the attack performance on frequency-based models. FPBA gets an average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than SSA, MIFGSM and IFGSM respectively.
Finally, FPBA has nearly top-2 results in all 48 cases (datasets vs. victim models vs. target models vs. attack setting), which shows consistent attack performance across different attack settings (white-box vs. black-box), tasks (AIGI detection vs. Deepfake detection) and models (CNN-based vs. Vit-based vs. Frequency-based detectors). We leave the detailed robustness analysis across models to Sec. 4.4.1.
Figure 5. Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGan datasets) for frequency-based models. The frequency components of frequency-based models are highly sparse in comparison with spatial-based models.
Table 2. The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of ResNet-50, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of ResNet, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of ResNet, DCTA and ViT. 4 means only using ResNet as surrogate model.
Attack ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave
ENS1 55.9 45.4 69.9 76.2 16.8 42.9 23.3 43.7 34.4
SVRE1 63.7 64.5 75.5 72.9 10.3 30.4 28.6 64.2 39.6
ENS2 57.3 58.6 54.1 92.8 19.5 38.5 77.9 59.7 46.0
SVRE2 77.7 77.9 76.4 97.8 13.9 33.6 70.1 78.1 56.0
ENS3 9.7 6.3 4.8 10.1 37.0 72.9 20.6 5.5 12.7
SVRE3 22.6 12.9 8.1 13.2 29.4 87.0 24.0 9.6 14.6
FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
4.2.1. Comparsion with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods, including ENSEMBLE  (Dong et al., 2018) and SVRE (Xiong et al., 2022). Since there are three types of models for AIGI detection (CNN-based vs. Vit-based vs. Frequency-based detectors), we hence conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (ResNet-50, MobileNet, EfficientNet); (2) taking CNNs and Vits as ensemble surrogates (ResNet-50, MobileNet, Vit); (3) taking CNNs, Vits and frequency-based detectors as ensemble surrogates (ResNet-50, DCTA, Vit). Although we can also use more than one architecture for our method, we only use ResNet-50 as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Tab. 2. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across Vit-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on ResNet still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by ResNet-50 with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of ResNet-50, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
Table 3. Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot (Wang et al., 2020).
Attack GramNet LGrad LNP UnivFD Ave
IFGSM 39.4 7.5 37.0 14.6 24.6
MIFGSM 9.65 6.6 38.5 24.2 19.7
PGD 67.7 50.2 47.4 11.5 44.2
S2I 49.1 49.1 28.3 8.3 33.7
FPBA 87.8 50.2 39.5 16.6 48.5
4.3. The Evaluation on Diverse AIGI Detections
4.3.1. Attack SOTA AIGI Detectors
In this section, we examine the robustness of state-of-the-art (SOTA) AIGI detectors against adversarial attacks and present the results in Tab. 3. Even though these detectors can achieve SOTA detection performance on benign data, they still suffer from a serious adversarial threat and demonstrate our method is the strongest threat to these SOTA detectors. Unlike common detectors trained on the fake images for the task of binary classification, UnivFD (Ojha et al., 2023) extract fixed features via CLIP (Radford et al., 2021) and use nearest neighbor or linear probing for classification. The obvious diversity of the surrogate model(CNNSpot) and victim model(UnivFD) leads to a limited adversarial transferability on each other. Nevertheless, our method still achieves the best results on UnivFD compared to other attacks.
Table 4. Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
Detector Acc Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
All 60.5 100.0 99.9 50.8 99.7 50.0 76.8
Real 100.0 100.0 100.0 100.0 100.0 99.9 100.0
Swin-ViT Fake 21.0 100.0 99.9 1.5 99.3 0.1 53.6
All 62.0 99.6 99.5 50.3 97.9 49.8 76.5
Real 99.4 99.6 99.6 99.5 99.6 99.4 99.5
ResNet Fake 24.7 99.6 99.4 1.2 96.1 0.2 53.5
All 58.4 99.5 99.3 50.2 96.0 50.2 75.6
Real 99.5 99.6 99.5 99.5 99.5 99.3 99.5
ViT Fake 17.4 99.3 99.1 0.9 92.4 1.1 51.7
Table 5. The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-Vit.
Victim Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Swin-ViT 97.6 97.4 97.7 96.2 96.9 97.1 97.1
ResNet 22.9 44.2 44.8 5.0 44.9 6.2 28.0
ViT 19.9 14.5 16.6 3.3 20.1 2.1 12.7
Table 6. The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-Vit.
Victim ASR Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
Real 94.9 96.4 96.2 96.0 96.0 97.0 95.9
Swin-ViT Fake 96.5 98.3 98.7 100.0 97.9 100.0 98.3
Real 3.0 3.4 2.9 2.0 3.6 4.6 3.2
ResNet Fake 100.0 85.1 86.9 100.0 86.3 100.0 93.0
Real 0.1 0.0 0.3 0.3 0.4 0.4 0.3
ViT Fake 96.1 29.0 32.9 100.0 39.8 100.0 66.3
4.3.2. Attack Cross-Generator Image Detection
One important real-world detection problem is cross-generator image detection, i.e. identifying fake images generated by unseen generative models. We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection. We train the detectors on images generated by SD v1.4 (Rombach et al., 2022) and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney (Midjourney, 2022), SD V1.4 (Rombach et al., 2022), SD V1.5 (Rombach et al., 2022), ADM (Dhariwal and Nichol, 2021), Wukong (Wukong, 2022) and BigGAN (Brock et al., 2018). Because Swin-Vit achieves the SOTA results on different subsets (Zhu et al., 2024), we use it as the surrogate model.
The benign accuracy and attack performance on unseen source data are reported in Tab. 4 and Tab. 5 respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Tab. 4,Tab. 6), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
4.3.3. Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training (Madry et al., 2017) for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically (Wang et al., 2020) uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from (Wang et al., 2020; Zhu et al., 2024) to utilize Gaussian blurring and JPEG compression as a defense. As shown in Tab. 7, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
Table 7. Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
Swin-ViT
Blur+JPEG(0.1) Blur+JPEG(0.5)
IFGSM 98.8 86.6(-12.2)
MIFGSM 98.3 92.7(-5.6)
PGD 99.3 87.3(-12)
SSA 100.0 80.9(-19.1)
FPBA 100.0 99.2(-0.5)
4.4. Additional Performance Analysis
4.4.1. The Phenomenon of Gradient Masking in AIGI Detectors
In Tab. 1, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate (Dong et al., 2018). It indicates the possibility of gradient masking (Athalye et al., 2018). To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by ResNet-50 on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 6 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.
Most gradient components of adversarial examples generated by IFGSM (Fig. 6(a)) and MIFGSM (Fig. 6 (c)) tend to stabilize around zero, indicating the vanishing gradient leading to a limited transferability to IFGSM and MIFGSM. Because the only difference between IFGSM and PGD is that PGD randomly chooses the starting point within the $l_{\infty}$ constraint, we apply random initialization to MIFGSM and find the value of gradient component increase (Fig. 6 (d)), w.r.t. the average success rate increasing from 40.4% to 58.6% (still lower 10.6% than ours). This analysis formally demonstrates the phenomenon of gradient masking in AIGI detectors. Therefore, we advocate for future work for attacks on AIGI detectors to employ randomized-based strategies to circumvent the effect of gradient masking. Our proposed method conducts the spectrum transformation in the frequency domain and hence is also effective for gradient masking.
Figure 6. The gradient components of ResNet on adversarial examples generated by different attack methods. (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
4.4.2. Qualitative Analysis
Table 8. The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
[…]
PGD 30.00 33.51 0.88
FPBA(Ours) 16.08 36.26 0.94
Figure 7. Visual comparison with PGD. (a): the original image generated by diffusion models. (b): adversarial examples crafted by PGD. (c): adversarial examples crafted by FPBA. The image quality of the adversarial example crafted by our method is much closer to the original image.
In this study, we evaluated the image quality of the generated adversarial samples using PSNR and SSIM scores as metrics. We present the quantitative results of MSE, PSNR(db) and SSIM in Tab. 8. The image quality of our method is higher than PGD. As illustrated in Fig. 7, FPBA can generate more natural-looking adversarial examples in comparison with PGD.
4.4.3. Ablation Study
Table 9. Ablation Study on ProGAN dataset. The adversarial samples are crafted from ResNet50.
Attack DenNet MobNet Spec ViT
Spatial 71.5 34.0 27.1 11.0
Frequency 91.0 64.7 36.7 16.9
Spatial-frequency 96.2 66.2 47.9 21.0
We conduct an ablation study in Tab. 9 to investigate the impact of computing gradient in different domains. In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability. More ablation studies can be found in the supplementary material.
5. Conclusion
In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks. For the task of AIGI detection, we propose a novel frequency-based post-train Bayesian attack, which explores the vulnerable regions in the frequency domain in a Bayesian manner. More broadly, our proposed post-train Bayesian attack strategy extends the Bayesian attack family. Extensive experiments have been conducted across models, generators and defense methods under both white-box and black-box settings. The following conclusions can be drawn from our investigation.
  • The current SOTA AIGI detectors are vulnerable to adversarial attacks, while the widely-used defense mechanism adversarial training cannot be directly applied to the AIGI detectors, which brings great challenges to build reliable and robust detectors.
  • AIGI detectors rely on gradient masking to mitigate the transferability of gradient-based attacks, but the effect of gradient masking can be easily circumvented by a simple random initialization strategy.
  • The attack success rate is positively associated with the detection accuracy. Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
These findings indicate that building robust AIGI detectors is still an open problem. We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to the robustness of AIGI detectors.

### v3

@@
 Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks
-Abstract.
+Abstract
 Recent advancements in image synthesis, particularly with the advent of GAN and Diffusion models, have amplified public concerns regarding the dissemination of disinformation.
@@
 In this paper, we examine the vulnerability of state-of-the-art AIGI detectors against adversarial attack under white-box and black-box settings, which has been rarely investigated so far.
-For the task of AIGI detection, we propose a new attack containing two main parts.
+To this end, we propose a new method to attack AIGI detectors.
 First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution.
-Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models, e.g.
+Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous AIGI detectors, e.g.
 transferring adversarial examples across CNNs and ViTs.
@@
 The code will be shared upon acceptance.
-AI-generated Image Detection, Adversarial Examples
-ccs: Security and privacy Software and application securityccs: Security and privacy Human and societal aspects of security and privacyccs: Computing methodologies Computer vision
-1. Introduction
-The notable progress in generative models, such as GANs (Goodfellow et al., 2014a) and Diffusion models (Ho et al., 2020), is driving the flourishing development of the image synthesis domain.
+Index Terms:
+AI-generated Image Detection, Adversarial Examples.
+I Introduction
+Figure 1: A high-level illustration of our proposed method.
+The notable progress in generative models, such as GANs [1] and Diffusion models [2], is driving the flourishing development of the image synthesis domain.
 These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images.
@@
 However, alongside the benefits, the widespread availability of fake images raises concerns regarding the dissemination of misinformation and fake media news.
-Consequently, numerous detectors have been proposed(Durall et al., 2020; Chai et al., 2020; Corvi et al., 2023b) to identify AI-generated images (AIGI).
-Recent state-of-the-art detectors (Wang et al., 2020; Zhong et al., 2023) rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
+Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI).
+Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
 However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real.
-Despite several works (Carlini and Farid, 2020; Hussain et al., 2021; Neekhara et al., 2021; Hou et al., 2023; Jia et al., 2022a) exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored.
+Despite several works [5, 6, 7, 8, 9] exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored.
 Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs.
 These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
-Figure 1. Illustration of the proposed method.
 In this paper, we show that AIGI detectors are vulnerable to adversarial attacks.
-Considering that many works (Dzanic et al., 2020; Frank et al., 2020) have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
+Considering that many works [10, 11] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
 As illustrated in Fig.
 3, the frequency components that different detectors focus on significantly vary from each other.
-Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models.
-To further improve the adversarial transferability across heterogeneous models, e.g.
-transferring adversarial examples across Visual Transformers (ViTs) and Convolutional Neural Networks (CNNs), we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate.
+Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains.
+Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[12], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g.
+transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table I).
+To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate.
 In contrast to existing ensemble-based or Bayesian attacks, which involve retraining an ensemble of surrogate models, our post-train Bayesian strategy freezes the pre-trained surrogate and appends tiny extra Bayesian components behind the surrogate, avoiding a heavy memory footprint and speeding up the training process.
-As a result, we propose the first adversarial attack for general AIGI detection, to add adversarial perturbations in the frequency domain from a post-train Bayesian perspective.
+As a result, we propose a new transferable adversarial attack for general AIGI detection, to add adversarial perturbations in various frequency transformation domains from a post-train Bayesian perspective.
 We name our method Frequency-based Post-train Bayesian Attack, or FPBA.
@@
 The contributions of this work can be summarized as follows:
-  • To our best knowledge, we are the first to systemically investigate the adversarial vulnerability of state-of-the-art AIGI detectors.
-We extensively evaluate the adversarial robustness of traditionally trained models, defense models, and cross-generated detection within real-world scenarios.
+  • We systemically investigate the adversarial robustness of state-of-the-art AIGI detectors, including standard trained models, defense models, and cross-generated detection within real-world scenarios.
 We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
@@
 A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
-  • We conduct comprehensive experiments on eight spatial-based detectors with CNN-based, Vit-based and CLIP-based architectures, and two frequency-based detectors using AIGI and Deepfake datasets.
-The extensive results demonstrate our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
-2. Related Work
-AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs (Goodfellow et al., 2014a), VAEs (Kingma and Welling, 2014), and Diffusion models (Ho et al., 2020), has brought about substantial risks in the dissemination of disinformation.
-Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods to primarily focus on face forgeries.
+  • We conduct comprehensive experiments on 13 various AIGI detectors, including heterogeneous model architectures and various detection methods.
+Extensive results demonstrate that our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
+II Related Work
+AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs [1], VAEs [13], and Diffusion models [2], has brought about substantial risks in the dissemination of disinformation.
+Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods [14, 15, 16], to primarily focus on face forgeries.
 However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects.
@@
 The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances.
-Data-driven-based approaches (Durall et al., 2020; Chai et al., 2020; Wang et al., 2020; Corvi et al., 2023b) have demonstrated promising results on generated data identically distributed.
+Data-driven-based approaches [3, 17] have demonstrated promising results on generated data identically distributed.
 However, their performance significantly degraded when presented with previously unseen generated images.
-More recently, certain methods (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) focus on analyzing forgery traces in the frequency domain and learning relevant forgery features (Corvi et al., 2023a; Liu et al., 2023; Pontorno et al., 2024; Zhang et al., 2019) to enhance the generalization capability of the detection system.
-In pursuit of these objectives, some methods have been proposed to learn local forgery features (Ju et al., 2023; Lorenz et al., 2023) in the spatial domain.
-AIGI fingerprint representations (Jeong et al., 2022) are acquired through the learning of noise patterns.
-Notably, studies such as (Liu et al., 2022; Zhong et al., 2023) have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance.
-A sequence of detection methodologies (Tan et al., 2023; Liu et al., 2023; Cozzolino et al., 2023; Chang et al., 2023) has been progressively investigated, relying on the practice of freezing pre-trained models.
+More recently, certain methods [10, 11] focus on analyzing forgery traces in the frequency domain and learning relevant forgery features [18] to enhance the generalization capability of the detection system.
+In pursuit of these objectives, some methods have been proposed to learn local forgery features [19, 20] in the spatial domain.
+AIGI fingerprint representations [21] are acquired through the learning of noise patterns.
+Notably, studies such as [22, 4] have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance.
+A sequence of detection methodologies [23] has been progressively investigated, relying on the practice of freezing pre-trained models.
 These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers.
-For instance, UniFD (Ojha et al., 2023) explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection.
-Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error (Wang et al., 2023a; Ricker et al., 2024), adversarial teacher-student discrepancy-aware framework (Zhu et al., 2023a) or diffusion noise(Zhang and Xu, 2023).
+For instance, UniFD [24] explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection.
+Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error [25, 26], adversarial teacher-student discrepancy-aware framework [27] or diffusion noise[28].
 The existing models consistently demonstrate enhancement in detecting AI-generated images.
 However, their behaviors under adversarial attacks have not been thoroughly examined.
-Our work is complementary to existing research by making the first attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
+Our work is complementary to existing research by making the attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
 Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly.
-The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including face recognition (Jia et al., 2022b), autonomous driving (Zhu et al., 2023b; Xie et al., 2023) and human activity recognition (Diao et al., 2021).
+The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including image classification[29, 30], multimedia communication [31] and human activity recognition [32].
 Very recently, the adversarial attack against Deepfake detection has drawn attention.
 Hussain et al.
-(Hussain et al., 2021) employ gradient sign-based methods (Kurakin et al., 2018; Goodfellow et al., 2014b) to evaluate the adversarial robustness of video Deepfake detectors.
+[6] employ gradient sign-based methods [33, 34] to evaluate the adversarial robustness of video Deepfake detectors.
 Li et al.
-(Li et al., 2021) generate adversarial fake images along the latent face manifold.
+[35] generate adversarial fake images along the latent face manifold.
 Carlini et al.
-(Carlini and Farid, 2020) and Neekhara et al.
-(Neekhara et al., 2021) investigate the adversarial threat to Deepfake detection under the black-box setting.
+[5] and Neekhara et al.
+[7] investigate the adversarial threat to Deepfake detection under the black-box setting.
 Jia et al.
-(Jia et al., 2022a) propose to generate adversarial perturbations in the frequency domain.
+[9] propose to generate adversarial perturbations in the frequency domain.
 Hou et al.
-(Hou et al., 2023) narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
+[8] narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
+FakePolisher [36] and TraceEvader [37] improve the untraceability of deepfakes by removing artificial fingerprints.
 Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation.
@@
 In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
-3. Methodology
-Figure 2. The workflow of FPBA. We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
-3.1. Preliminaries
+III Methodology
+Figure 2: The workflow of FPBA.
+We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
+DCT and IDCT are discrete cosine transformation and inverse discrete cosine transformation respectively.
+III-A Preliminaries
 Let $\mathbf{x}$ and $y$ represent the original image and its corresponding label.
@@
 maximizing the classification loss:
-(1) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t.
-}\left\|\delta\right\|_{p}\leq\epsilon$
+$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t.
+}\left\|\delta\right\|_{p}\leq\epsilon,$ (1)
 where $L$ is the binary cross entropy loss in AI-generated image detection.
@@
 Eq.
-1 can be performed with iterative gradient-based methods, such as PGD (Madry et al., 2017) or I-FGSM (Kurakin et al., 2018):
-(2) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
-(3) $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$
-3.2. Frequency-based Analysis and Attacks
+1 can be performed with iterative gradient-based methods, such as PGD [38] or I-FGSM [33]:
+$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
+$=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}(\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta)).$ (2)
 […]
-Many AIGI detection approaches distinguish between real and fake images via subtle artifacts (Wang et al., 2020; Jeong et al., 2022; Zhong et al., 2023).
-While these subtle clues are invisible in the spatial domain, a series of works (Dzanic et al., 2020; Frank et al., 2020; Chandrasegaran et al., 2021) demonstrate that there are obvious differences between real and fake images in the frequency domain.
+III-B Frequency-based Analysis and Attacks
+Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 21, 4].
+While these subtle clues are invisible in the spatial domain, a series of works [10, 11] demonstrate that there are obvious differences between real and fake images in the frequency domain.
 This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective.
 To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain.
-To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map (Long et al., 2022) to visualize the sensitive components of real and fake images across different models:
-(4) $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})}$
+To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [39] to visualize the sensitive components of real and fake images across different models:
+$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
 […]
-(5) $\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$
-where $\Gamma(\cdot)$ denotes the random spectrum transformation (Long et al., 2022).
+$\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
+where $\Gamma(\cdot)$ denotes the random spectrum transformation [39].
 $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$.
@@
 We hence conduct adversarial attack in the frequency domain via spectrum transformation:
-(6) $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon$
-3.3. Exploring the Surrogate Posterior Space
+$\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon.$ (5)
+III-C Exploring the Surrogate Posterior Space
 Although tuning the spectrum transformation in Eq.
-6 can simulate different substitute models with a homogeneous architecture (Long et al., 2022), it shows limited transferability when applied to heterogeneous architectures, e.g.
+5 can simulate different substitute models with a homogeneous architecture [39], it shows limited transferability when applied to heterogeneous architectures, e.g.
 transferring adversarial examples across ViTs and CNNs.
@@
 $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,p(y\mid\Gamma(\tilde{\mathbf{x}}),\mathcal{D})$
-(7) $=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t.
-}\left\|\delta\right\|_{p}\leq\epsilon$
+$=$ $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}\,\mathbb{E}_{\Theta\sim p(\Theta\mid\mathcal{D})}p\left(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta\right),\text{ s.t.
+}\left\|\delta\right\|_{p}\leq\epsilon,$ (6)
 where $p(\Theta\mid\mathcal{D})\propto p(\mathcal{D}\mid\Theta)p(\Theta)$.
@@
 Attacking Bayesian Neural Networks (BNNs) rather than a single DNN allows for the output fusion from an ensemble of infinitely many DNNs with diverse predictions, thereby improving adversarial transferability.
-Figure 4. The architecture of the appended model. $\sigma$ means the sigmoid layer.
-3.3.1. Post-train Bayesian Strategy for Frequency-based Attack
+Figure 4: The architecture of the appended model. $\sigma$ means the sigmoid layer.
+III-C1 Post-train Bayesian Strategy
 However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors.
-First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs (Izmailov et al., 2021).
+First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs [40].
 Hence computing and sampling the posterior distribution is an intractable problem.
 Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space.
-Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets (Zhu et al., 2024; He et al., 2021).
+Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [12, 41].
 From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
@@
 The new logits can be computed via a skip connection:
-(8) $\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x})$
+$\operatorname{logits}=g_{\theta^{\prime}}(f_{\theta}(\mathbf{x}))+f_{\theta_{c}}(\mathbf{x}).$ (7)
 We choose to apply Bayesian Model Averaging to optimize the appended Bayesian model:
-(9) $\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)$
+$\mathbb{E}_{\theta^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta)}p\left(y\mid\mathbf{x},\Theta,\theta^{\prime}\right)\approx\frac{1}{K}\sum_{k=1}^{K}p\left(y\mid\mathbf{x},\Theta,\theta_{k}^{\prime}\right),$
+$\theta_{k}^{\prime}\sim p(\theta^{\prime}\mid\mathcal{D},\Theta),$ (8)
 where $K$ is the number of appended models.
@@
 Finally, the frequency-based post-train Bayesian attack can be conducted with iterative gradient-based methods:
-(10) $\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}$
-3.3.2. Inference on Bayesian Appended Models
-$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo(Springenberg et al., 2016) to sample appended model $\theta^{\prime}$ in each iteration:
-$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I})$
-(11) $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2}$
-where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically (Springenberg et al., 2016).
-3.4. Hybrid Adversarial Attack
-Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain (Wang et al., 2020; Jeong et al., 2022).
+$\tilde{\mathbf{x}}^{i+1}=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}),\Theta,\theta^{\prime}_{k})\}.$ (9)
+III-C2 Inference on Bayesian Appended Models
+$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[42] to sample appended model $\theta^{\prime}$ in each iteration:
+$\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I}),$
+$\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2},$ (10)
+where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [42].
+III-D Hybrid Adversarial Attack
+Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3, 21].
 We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains.
 Specifically, we define the hybrid attack as:
-(12) $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\}$
-(13) $g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i}$
-(14) $d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k})$
+$\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}-\alpha\cdot\operatorname{sign}\{\frac{1}{K}\sum_{k=1}^{K}(g^{i}_{k}+d^{i}_{k})\},$ (11)
+$g^{i}_{k}$ $=\frac{1}{N}\sum_{n=1}^{N}\nabla\,log\,p(y\mid\Gamma(\tilde{\mathbf{x}}^{i}_{n-1}),\Theta,\theta^{\prime}_{k}),\tilde{\mathbf{x}}^{i}_{0}=\tilde{\mathbf{x}}^{i},$ (12)
+$d^{i}_{k}$ $=\nabla\,log\,p(y\mid\tilde{\mathbf{x}}^{i},\Theta,\theta^{\prime}_{k}),$ (13)
 where $g^{i}_{k}$ and $d^{i}_{k}$ are the gradients computed in the frequency domain and spatial domain respectively.
@@
 2.
-Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models; ;
+Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
 Output: The adversarial example $\tilde{\mathbf{x}}$;
@@
              for t = 1 to $M_{\theta^{\prime}}$ do
-                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Sec. 3.3.2;
+                   Update $\theta^{\prime}_{k}$ with $\mathbf{h}_{\theta^{\prime}_{k}}$ via Section III-C2;
              end for
@@
       for k = 1 to $K$ do
-             Averaging frequency gradient $g_{k}$ using Eq. 13 ;
-             Calculating spatial gradient $d_{k}$ using Eq. 14 ;
+             Average frequency gradient $g_{k}$ using Eq. 12 ;
+             Calculate spatial gradient $d_{k}$ using Eq. 13 ;
        end for
@@
 Algorithm 1 Inference on FPBA
-4. Experimets
-4.1. Experimental Settings
+IV Experimets
+IV-A Experimental Settings
 Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection.
-Synthetic LSUN is a commonly used dataset proposed by CNNSpot (Wang et al., 2020), containing 720k real images from LSUN and 360k fake images generated by ProGAN (Karras et al., 2018).
-GenImage (Zhu et al., 2024) is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
-Following the protocol in (Zhu et al., 2024), we employ a subset of GenImage collecting 162k real images from Imagenet (Deng et al., 2009) and 162k Stable Diffusion(SD) V1.4 (Rombach et al., 2022) generated images for training.
-The images generated by the other generators are used for testing in Sec.
-4.3.2.
-To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by (Shamshad et al., 2023).
-The Deepfake dataset consists of 50k real face images from FFHQ (Karras et al., 2019) and 50k generated face images generated by StyleGAN2 (Karras et al., 2019).
+Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43].
+GenImage [12] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
+Following the protocol in [12], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training.
+The images generated by the other generators are used for testing in Section IV-C2.
+To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46].
+The Deepfake dataset consists of 50k real face images from FFHQ [47] and 50k generated face images generated by StyleGAN2 [47].
 After training, we collect only the correctly classified testing samples for attack in evaluation.
-Evaluated Models: We extensively evaluate the transferability of adversarial examples by eight different models across heterogeneous architectures under the spatial and frequency domain.
-We choose spatial-based classification models ResNet-50 (He et al., 2016) and MobileNet (Howard et al., 2017) as the source models.
-For target models, apart from the above models, we also use spatial-based models including EfficientNet (Tan and Le, 2019), Vision Transformer(Vit) (Dosovitskiy et al., 2020) and Swin-Transformer(Swin-Vit) (Liu et al., 2021), and frequency-based detectors DCTA (Frank et al., 2020) and Spec (Zhang et al., 2019).
-We report the accuracy of these detectors in the supplement.
-Besides, we also evaluate the robustness of state-of-the-art AIGI detection methods, including GramNet (Liu et al., 2020), LGrad (Tan et al., 2023), LNP (Liu et al., 2022) and UnivFD (Ojha et al., 2023).
-Compared Methods: We adopt gradient-based methods, I-FGSM (Kurakin et al., 2018), PGD (Madry et al., 2017), MI (Dong et al., 2018) and Frequency-based methods S2(Long et al., 2022).
-We also consider ensemble-based methods Ensemble (Dong et al., 2018) and SVRE (Xiong et al., 2022).
-For a fair comparison, we ran 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all attacks.
-Implementation Details: We choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001.
-We follow the same data augmentation strategy in CNNSpot (Wang et al., 2020).
-Before cropping, images are blurred with $\sigma$ Uniform[0, 3] with 10% probability, and JPEG-ed with 10% probability.
-We transformed images to 224 pixels and apply ImageNet normalization.
-For post-train Bayesian optimization, we follow the default setting in (Wang et al., 2023b).
+Evaluated Models: We extensively evaluate the transferability of adversarial examples by 13 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods.
+For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [48], EfficientNet [49] and DenseNet [50], and ViT-based detectors Vision Transformer(ViT) [51] and Swin-Transformer(Swin-ViT) [52].
+For evaluating on various detection methods, we use frequency-based detectors DCTA [11] and Spec [18], gradient-based detectors LGrad [23], CLIP-based detectors UnivFD [24], and diffusion-based detectors DNF [28].
+In addition, LNP [22] extracts the noise pattern of images and GramNet [53] learns the global texture representation.
+We also consider them as victim models.
+Compared Methods: We adopt gradient-based methods, I-FGSM [33], PGD [38], MI [54] and frequency-based methods S2[39], SSAH [55] and ensemble-based methods ENS [54] and SVRE [56].
+Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks Fakepolisher [36] and TraceEvader [37] as baselines.
+Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings.
+For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
+Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3].
+Specificaly, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001.
+For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness.
+Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability.
+We transformed images to 224 pixels on LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3].
+On the FFHQ (StyleGAN2) dataset we resized images to 224 pixels to ensure the integrity of the real/fake faces.
+Subsequently, we apply ImageNet normalization across three datasets.
+For other AIGI detectors, we use the pre-trained model from their official code.
+For post-train Bayesian optimization, we follow the default setting in [57].
 Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate.
 Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$.
-For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following (Long et al., 2022).
+For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [39].
 All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
-Table 1.
-The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN, GenImage subset and Synthetic FFHQ datasets.
+TABLE I: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset.
 “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
-Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
+We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
+Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
 IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
 MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
-ResNet PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
-S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
-FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
+PGD 78.3 77.6 49.4 73.0 25.6 47.6 41.0 70.2 54.9
+CNNSpot S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
+SSAH 97.8 2.5 1.3 3.0 0.8 1.0 1.0 0.7 1.4
+FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
 IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
-MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
-MobileNet PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
-S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
+MIFGSM 23.8 23.9 28.6 75.4 19.6 42.0 15.5 24.8 25.5
+PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
+MobileNet S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
+SSAH 0.6 1.0 1.5 99.2 0.5 0.8 1.0 1.1 0.9
 LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
@@
 MIFGSM 78.3 75.3 30.1 35.8 27.8 28.3 15.0 28.2 34.4
-ResNet-50 PGD 99.3 88.0 48.5 52.2 48.6 49.9 15.4 48.1 50.1
-S2I 98.0 91.0 66.8 64.7 36.7 50.1 16.9 53.0 54.2
+PGD 99.3 88.0 48.5 52.2 48.6 49.9 15.4 48.1 50.1
+CNNSpot S2I 98.0 91.0 66.8 64.7 36.7 50.1 16.9 53.0 54.2
+SSAH 94.8 5.4 0.5 1.1 0.6 3.6 0.7 0.3 1.7
 FPBA(Ours) 100 96.6 59.4 67.3 47.9 49.8 21.1 51.2 56.2
@@
 MIFGSM 9.7 9.2 8.9 59.0 9.5 9.0 5.9 8.8 8.7
-MobileNet PGD 42.3 39.9 36.4 97.1 48.4 49.8 14.2 40.0 38.7
-S2I 35.6 33.1 26.8 78.7 24.1 43.4 6.2 28.0 28.2
+PGD 42.3 39.9 36.4 97.1 48.4 49.8 14.2 40.0 38.7
+MobileNet S2I 35.6 33.1 26.8 78.7 24.1 43.4 6.2 28.0 28.2
+SSAH 1.1 2.0 0.6 96.7 0.5 1.9 1.2 2.2 1.4
 GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
-IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
-MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
-ResNet PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
-S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
-FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
-IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
-MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
-MobileNet PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
-S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
-FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
-4.2. The Evaluation on Normally Trained Models
-We show the attack performance against all models on all datasets in Tab.
-1.
+IV-B Attack on Spatial-based and Frequency-based Detectors
+We report the attack performance against spatial-based and frequency-based detectors in Table I.
 Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods.
-Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.8%, while S2I, PGD, IFGSM and MIFGSM only have 95.6%, 95.4%, 65.4% and 65.5%.
-Under the black-box setting, FPBA still achieves the highest average transfer success rate of 55.3%, surpassing MIFGSM, PGD and S2I by a margin of 27.7%, 7.6% and 8.4% respectively.
-Next, We find that the attack results for FPBA are usually not the best on frequency-based models.
-To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq.
-4 in Fig.
-5.
-Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig.
-3), the frequency components focused on in frequency-based models are very sparse.
-Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain.
-Nevertheless, our method still has the best average success rate over all models.
-In addition, we combine spatial-based gradient and frequency-based to further improve the attack performance on frequency-based models.
-FPBA gets an average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than SSA, MIFGSM and IFGSM respectively.
-Finally, FPBA has nearly top-2 results in all 48 cases (datasets vs.
-victim models vs.
-target models vs.
-attack setting), which shows consistent attack performance across different attack settings (white-box vs.
-black-box), tasks (AIGI detection vs.
-Deepfake detection) and models (CNN-based vs.
-Vit-based vs.
-Frequency-based detectors).
-We leave the detailed robustness analysis across models to Sec.
-4.4.1.
-Figure 5.
-Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGan datasets) for frequency-based models.
+Specifically, FPBA gets an average white-box success rate across different datasets and models as high as 99.6%, while S2I, PGD, IFGSM and MIFGSM only have 93.1%, 93.1%, 66.2% and 66.2%.
+Although SSAH has a relatively high white-box success rate (97.7%), it fails to transfer the adversarial examples to black-box models.
+Under the black-box setting, FPBA still achieves the highest average transfer success rate of 50.5%, surpassing the IFGSM, MIFGSM, PGD, S2I and SSAH by a big margin of 25.7%, 23.3%, 8.4%, 11.9% and 49.1%.
+Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
+TABLE II: The attack success rate(%) compared with attacks agaisnt Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.
+Datasets Attack CNNSpot DenNet EffNet MobNet ViT Swin Ave.
+FakePolisher 94.5 98.3 71.6 88.6 15.1 99.5 77.9
+LSUN TraceEvader 0.6 1.3 4.5 4.8 8.10 11.4 1.9
+(ProGAN) FPBA 100.0 100.0 98.2 99.5 38.6 99.9 89.4
+FakePolisher 93.5 93.2 12.2 95.5 33.5 97.6 70.9
+GenImage TraceEvader 46.1 34.3 9.4 37.0 15.2 44.3 31.1
+(SD) FPBA 100.0 99.7 97.7 98.9 42.1 97.9 89.4
+Figure 5: Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGAN datasets) for frequency-based models.
 The frequency components of frequency-based models are highly sparse in comparison with spatial-based models.
-Table 2.
-The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets.
+The color value represents the absolute gradient value of the model loss function after max-min normalization.
+IV-B1 Comparison with Attacks against Deepfake Detectors
+TABLE III: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets.
 1 means setting (1), i.e.
-taking an ensemble of ResNet-50, MobileNet, and EfficientNet; 2 means setting (2), i.e.
-taking an ensemble of ResNet, MobileNet and ViT; 3 means setting (3), i.e.
-taking an ensemble of ResNet, DCTA and ViT.
-4 means only using ResNet as surrogate model.
-Attack ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave
+taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e.
+taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e.
+taking an ensemble of CNNSpot, DCTA and ViT.
+4 means only using CNNSpot as surrogate model.
+Attack CNNSpot DenNet EffNet MobNet Spec DCTA ViT Swin Ave
 ENS1 55.9 45.4 69.9 76.2 16.8 42.9 23.3 43.7 34.4
@@
 FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
-4.2.1. Comparsion with Ensemble-based Attacks
-Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods, including ENSEMBLE  (Dong et al., 2018) and SVRE (Xiong et al., 2022).
-Since there are three types of models for AIGI detection (CNN-based vs.
-Vit-based vs.
-Frequency-based detectors), we hence conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (ResNet-50, MobileNet, EfficientNet); (2) taking CNNs and Vits as ensemble surrogates (ResNet-50, MobileNet, Vit); (3) taking CNNs, Vits and frequency-based detectors as ensemble surrogates (ResNet-50, DCTA, Vit).
-Although we can also use more than one architecture for our method, we only use ResNet-50 as the surrogate architecture to verify the universal transferability across heterogeneous models.
-We report the results in Tab.
-2.
-First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across Vit-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors.
-We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate.
-Second, FPBA solely on ResNet still achieves competitive results in comparison with ensemble methods.
-Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases.
-For instance, the adversarial example only generated by ResNet-50 with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of ResNet-50, EfficientNet and MobileNet.
-This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
-Table 3. Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot (Wang et al., 2020).
-Attack GramNet LGrad LNP UnivFD Ave
-IFGSM 39.4 7.5 37.0 14.6 24.6
-MIFGSM 9.65 6.6 38.5 24.2 19.7
-PGD 67.7 50.2 47.4 11.5 44.2
-S2I 49.1 49.1 28.3 8.3 33.7
-FPBA 87.8 50.2 39.5 16.6 48.5
-4.3. The Evaluation on Diverse AIGI Detections
-4.3.1. Attack SOTA AIGI Detectors
-In this section, we examine the robustness of state-of-the-art (SOTA) AIGI detectors against adversarial attacks and present the results in Tab.
-3.
-Even though these detectors can achieve SOTA detection performance on benign data, they still suffer from a serious adversarial threat and demonstrate our method is the strongest threat to these SOTA detectors.
-Unlike common detectors trained on the fake images for the task of binary classification, UnivFD (Ojha et al., 2023) extract fixed features via CLIP (Radford et al., 2021) and use nearest neighbor or linear probing for classification.
-The obvious diversity of the surrogate model(CNNSpot) and victim model(UnivFD) leads to a limited adversarial transferability on each other.
-Nevertheless, our method still achieves the best results on UnivFD compared to other attacks.
-Table 4. Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
+TABLE IV: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
+Attack GramNet LGrad LNP UnivFD DNF Ave
+MIFGSM 9.7 6.6 38.5 24.2 46.7 25.1
+PGD 67.7 50.2 47.4 11.5 48.3 45.0
+S2I 49.1 49.1 28.3 8.3 43.0 35.6
+FPBA 87.8 50.2 39.5 16.6 49.8 48.9
+TABLE V: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
 Detector Acc Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
@@
 Real 99.4 99.6 99.6 99.5 99.6 99.4 99.5
-ResNet Fake 24.7 99.6 99.4 1.2 96.1 0.2 53.5
+CNNSpot Fake 24.7 99.6 99.4 1.2 96.1 0.2 53.5
 All 58.4 99.5 99.3 50.2 96.0 50.2 75.6
@@
 ViT Fake 17.4 99.3 99.1 0.9 92.4 1.1 51.7
-Table 5. The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-Vit.
+Attacks against Deepfake detectors are similar to our attacks.
+However, Deepfake mainly manipulating or synthesizing faces, often crafted by GANs.
+In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models.
+The content includes humans, non-human entities, abstract art, landscapes, and more.
+These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors.
+For example, techniques like searching for adversarial points on the face manifold [35] or removing facial forgery traces [36, 37] may not be applicable for AIGI detectors.
+To demonstrate this, we compare with FakePolisher [36] and TraceEvader [37], which are both the SOTA attacks for Deepfake detection.
+Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images.
+As shown in Table II, FPBA significantly outperforms FakePolisher and TraceEvader on both AIGI datasets.
+Although FakePolisher performs better than TraceEvader, but the former largely reduces the image quality, which is very visible and raises suspicion (Fig.
+7).
+We suspect this is because FakePolisher has only learned the face manifold, limiting its ability to reconstruct fine-grained details in AIGI images.
+IV-B2 Comparison with Ensemble-based Attacks
+Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [54] and SVRE [56], i.e.
+utilizing an ensemble of surrogate models to generate adversarial examples.
+To examine the impact of different surrogate combinations(CNN-based vs.
+Vit-based vs.
+Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT).
+Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models.
+We report the results in Table III.
+First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors.
+We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate.
+Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods.
+Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases.
+For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet.
+This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
+IV-B3 Evaluation on Frequency-based Detectors
+Although FPBA has the best average success rate, we find that the attack results are usually not the best on frequency-based models.
+To further investigate the reason, we plot the spectrum saliency map of frequency-based detectors using Eq.
+3 in Fig.
+5.
+Unlike spatial-based detectors, which rely on numerous frequency components to make decisions (see Fig.
+3), the frequency components focused on in frequency-based models are very sparse.
+Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain.
+Nevertheless, FPBA still gets an top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM respectively.
+TABLE VI: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
 Victim Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
 Swin-ViT 97.6 97.4 97.7 96.2 96.9 97.1 97.1
-ResNet 22.9 44.2 44.8 5.0 44.9 6.2 28.0
+CNNSpot 22.9 44.2 44.8 5.0 44.9 6.2 28.0
 ViT 19.9 14.5 16.6 3.3 20.1 2.1 12.7
-Table 6. The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-Vit.
+TABLE VII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
 Victim ASR Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
@@
 Real 3.0 3.4 2.9 2.0 3.6 4.6 3.2
-ResNet Fake 100.0 85.1 86.9 100.0 86.3 100.0 93.0
+CNNSpot Fake 100.0 85.1 86.9 100.0 86.3 100.0 93.0
 Real 0.1 0.0 0.3 0.3 0.4 0.4 0.3
 ViT Fake 96.1 29.0 32.9 100.0 39.8 100.0 66.3
-4.3.2. Attack Cross-Generator Image Detection
+IV-C Evaluation on Diverse Detection Strategies
+IV-C1 Attack Various SOTA Detection Methods
+Except for detecting in spatial and frequency domian using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts.
+Specifically, GramNet [53] extracts the global texture representation; UnivFD [24] captures semantic feature using pre-trained CLIP [58]; LNP [22] extracts the noise pattern of images; LGrad [23] extract gradient information through a pre-trained model and DNF [28] leverages diffusion noise features, extracting them through an inverse diffusion process [2].
+These significant differences among various detection methods present challenges for the transferability of adversarial examples.
+To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table IV.
+Overall, the transfer success rate is relatively low compared to the results in Table I, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other.
+Nevertheless, FPBA still give the best adversarial transferability.
+IV-C2 Attack Cross-Generator Image Detection
 One important real-world detection problem is cross-generator image detection, i.e.
@@
 We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection.
-We train the detectors on images generated by SD v1.4 (Rombach et al., 2022) and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney (Midjourney, 2022), SD V1.4 (Rombach et al., 2022), SD V1.5 (Rombach et al., 2022), ADM (Dhariwal and Nichol, 2021), Wukong (Wukong, 2022) and BigGAN (Brock et al., 2018).
-Because Swin-Vit achieves the SOTA results on different subsets (Zhu et al., 2024), we use it as the surrogate model.
-The benign accuracy and attack performance on unseen source data are reported in Tab.
-4 and Tab.
-5 respectively.
+We train the detectors on images generated by SD v1.4 [45] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [59], SD V1.4 [45], SD V1.5 [45], ADM [60], Wukong [61] and BigGAN [62].
+Because Swin-ViT achieves the SOTA results on different subsets [12], we use it as the surrogate model.
+The benign accuracy and attack performance on unseen source data are reported in Table V and Table VI respectively.
 First, the attack under the white-box setting achieves almost 100% success rate.
@@
 In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them.
-By looking closely at the accuracy/ASR on real images and fake images(Tab.
-4,Tab.
-6), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack.
+By looking closely at the accuracy/ASR on real images and fake images(Table V,Table VII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack.
 Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
-4.3.3. Attack Defense Models
-Furthermore, we evaluate FPBA on attacking defense models.
-We first employ adversarial training (Madry et al., 2017) for evaluation, which is the most popular defense baseline.
-However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection.
-Considering that AIGI detection typically (Wang et al., 2020) uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from (Wang et al., 2020; Zhu et al., 2024) to utilize Gaussian blurring and JPEG compression as a defense.
-As shown in Tab.
-7, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
-Table 7.
-Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset.
+IV-D Attack Performance against Compression in Real Scenarios
+Figure 6: Attack performace on ResNet, Efficient, MobileNet and Swin against JPEG compression on LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
+In real-world scenarios, images are often compressed during propagation, which may also distort adversarial noise patterns.
+Therefore, ensuring the effectiveness of adversarial attacks on compressed images is crucial for simulating real-world scenarios.
+To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images.
+We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer.
+All adversarial samples are crafted from ResNet-50.
+The results are reported in Fig.
+6.
+For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance.
+For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases.
+However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
+TABLE VIII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset.
 Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively.
@@
 PGD 99.3 87.3(-12)
-SSA 100.0 80.9(-19.1)
-FPBA 100.0 99.2(-0.5)
-4.4. Additional Performance Analysis
-4.4.1. The Phenomenon of Gradient Masking in AIGI Detectors
-In Tab.
-1, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate (Dong et al., 2018).
-It indicates the possibility of gradient masking (Athalye et al., 2018).
+S2I 100.0 80.9(-19.1)
+FPBA 100.0 99.2(-0.8)
+IV-D1 Attack Defense Models
+Furthermore, we evaluate FPBA on attacking defense models.
+We first employ adversarial training [38] for evaluation, which is the most popular defense baseline.
+However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection.
+Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 12] to utilize Gaussian blurring and JPEG compression as a defense.
+As shown in Table VIII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
+TABLE IX: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets.
+“Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
+We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
+Surrogate Model Attack Methods CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Average
+IFGSM 67.5 67.5 55.3 67.5 28.9 32.5 16.2 17.5 40.8
+MIFGSM 67.5 67.5 54.9 67.6 28.8 32.6 20.7 17.9 41.4
+PGD 100 100 88.6 99.9 49.9 50 40.7 50.0 68.4
+CNNSpot S2I 100 100 77.7 90.1 46.7 52.9 37.6 55.8 65.8
+SSAH 99.0 54.5 44.1 41.0 2.6 7.2 1.0 4.2 22.1
+FPBA(Ours) 100 100 86.8 100 49.4 49.6 43.1 50.9 68.5
+IFGSM 10.7 12.7 23.8 60.4 24.5 28.4 7.3 10.2 16.8
+MIFGSM 10.3 12.6 23.2 60.4 25.4 28.7 9.1 10.2 17.1
+PGD 49.8 50.2 57.5 100 49.7 49.8 38.8 49.8 49.4
+MobileNet S2I 82.8 86.7 78.5 100 51.4 51.8 28.9 49.7 61.4
+SSAH 1.9 36.4 43.7 98.5 2.0 6.2 0.8 2.4 13.3
+FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
+IV-E Visual Quality Analysis
+TABLE X: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
+[…]
+PGD 30.00 33.51 0.88
+FakePolisher 34.63 32.92 0.88
+FPBA(Ours) 16.08 36.26 0.94
+Figure 7: Visual comparison with PGD.
+(a): the original image generated by diffusion models.
+(b): adversarial examples crafted by PGD.
+(c): the original image generated by FakePolisher.
+(d): adversarial examples crafted by FPBA.
+The image quality of the adversarial example crafted by our method is much closer to the original image.
+To demonstrate the superior image quality achieved by our method, we conduct both qualitative and quantitative assessments.
+We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig.
+7.
+Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in.
+In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images.
+In contrast, FPBA generate more natural-looking adversarial examples.
+Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM.
+As reported in Table X, FPBA outperforms other baselines across all quality assessment by a large margin.
+This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
+IV-F Additional Performance Analysis
+IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
+In Table I, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [54].
+It indicates the possibility of gradient masking [63].
 To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties.
-We take the adversarial examples crafted by ResNet-50 on ProGan as an example.
+We take the adversarial examples crafted by CNNSpot on ProGan as an example.
 For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients.
@@
 Most gradient components of adversarial examples generated by IFGSM (Fig.
-6(a)) and MIFGSM (Fig.
+8(a)) and MIFGSM (Fig.
 8 (c)) tend to stabilize around zero, indicating the vanishing gradient leading to a limited transferability to IFGSM and MIFGSM.
@@
 Our proposed method conducts the spectrum transformation in the frequency domain and hence is also effective for gradient masking.
-Figure 6.
-The gradient components of ResNet on adversarial examples generated by different attack methods.
+Figure 8: The gradient components of CNNSpot on adversarial examples generated by different attack methods.
 (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
-4.4.2. Qualitative Analysis
-Table 8. The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
-[…]
-PGD 30.00 33.51 0.88
-FPBA(Ours) 16.08 36.26 0.94
-Figure 7.
-Visual comparison with PGD.
-(a): the original image generated by diffusion models.
-(b): adversarial examples crafted by PGD.
-(c): adversarial examples crafted by FPBA.
-The image quality of the adversarial example crafted by our method is much closer to the original image.
-In this study, we evaluated the image quality of the generated adversarial samples using PSNR and SSIM scores as metrics.
-We present the quantitative results of MSE, PSNR(db) and SSIM in Tab.
-8.
-The image quality of our method is higher than PGD.
-As illustrated in Fig.
-7, FPBA can generate more natural-looking adversarial examples in comparison with PGD.
-4.4.3. Ablation Study
-Table 9. Ablation Study on ProGAN dataset. The adversarial samples are crafted from ResNet50.
-Attack DenNet MobNet Spec ViT
+IV-F2 Evaluation on Deepfake Datasets
+As we mention before, there is a big difference between AIGI detection and Deepfake detection.
+But we empirically found that FPBA is also effective for face forgery detection.
+To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [47].
+We report the results in Table IX.
+Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance.
+For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
+TABLE XI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
+Attack DenseNet MobileNet Spec ViT
 Spatial 71.5 34.0 27.1 11.0
@@
 Spatial-frequency 96.2 66.2 47.9 21.0
-We conduct an ablation study in Tab.
-9 to investigate the impact of computing gradient in different domains.
+TABLE XII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations.
+“Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
+$N$ ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
+0 64.5 64.0 37.3 60.6 17.5 48.4 40.5 59.7 48.8
+5 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
+10 99.6 99.1 78.6 97.5 19.9 48.0 52.4 96.5 70.3
+15 99.4 99.0 78.2 96.6 20.2 48.3 53.0 96.2 70.2
+20 99.2 98.7 77.5 96.7 20.7 48.6 51.1 95.6 69.8
+IV-G Ablation Study
+Spatial and Frequency Attack
+We conduct an ablation study in Table XI to investigate the impact of computing gradient in different domains.
 In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability.
 More ablation studies can be found in the supplementary material.
-5. Conclusion
+Number ($N$) of Spectrum Transformations
+In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models.
+The adversarial examples were crafted from ResNet-50 in LSUN(ProGAN) dataset.
+With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper.
+As shown in Table XII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate.
+Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation.
+Considering the trade-off between attack performance and computation cost, we set $N=5$ by default.
+V Conclusion
 In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks.
@@
 Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
-These findings indicate that building robust AIGI detectors is still an open problem.
-We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to the robustness of AIGI detectors.
+These findings indicate that building an adversarial robust AIGI detector is still an open problem.
+We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to this domain.
+Acknowledgments
+This work was supported in part by the NSF China (No. 62302139) and FRFCU-HFUT (JZ2023HGTA0202, JZ2023HGQA0101).

### v4

@@
 However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real.
-Despite several works [5, 6, 7, 8, 9] exploring adversarial attacks for GAN-based face forgery detection, we extend this security issue to encompass broader AIGI detection, which has been largely ignored.
-Unlike face forgery concentrates solely on manipulating or synthesizing faces, often crafted by GANs, AI-generated images span a broader spectrum, generated by a variety of advanced diffusion models and GANs.
-These images encompass a richer array of semantic content, thus greatly enhancing their potential for disseminating disinformation.
+Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging.
+Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [5, 6], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts.
+This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints.
+Consequently, transferring adversarial patterns across such diverse representations is far less straightforward.
+Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [7, 8, 9, 10, 11].
+These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces.
+In contrast, AIGI detectors are designed to identify any AI-generated content produced by a wide range of generative techniques, including GANs, diffusion models, and autoregressive models.
+The content extends beyond human faces to encompass non-human entities, landscapes, abstract art, and more.
+These differences make designing attacks for AIGI detection more challenging than Deepfake detection.
 In this paper, we show that AIGI detectors are vulnerable to adversarial attacks.
@@
 Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[14], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g.
-transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table I).
+transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II).
 To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate.
@@
 1.
-The contributions of this work can be summarized as follows:
-  • We systemically investigate the adversarial robustness of state-of-the-art AIGI detectors, including standard trained models, defense models, and cross-generated detection within real-world scenarios.
-We demonstrate adversarial examples are truly a real threat to AIGI detections, and reveal the phenomenon of gradient masking phenomena in AIGI detectors.
-  • We propose a novel attack against AIGI detection by exploring the vulnerable frequency region in a Bayesian manner.
-A new post-train Bayesian strategy for attack is proposed to explore the full posterior distribution over the surrogate, without the need for re-training the substitute model.
-  • We conduct comprehensive experiments on 13 various AIGI detectors, including heterogeneous model architectures and various detection methods.
-Extensive results demonstrate that our proposed method achieves the highest average attack success rate under white-box and black-box settings, outperforming baseline methods by a big margin.
+The contributions of this work can be summarized as follows: (1) We systematically assess the adversarial robustness of state-of-the-art AIGI detectors, revealing both real-world threats and gradient masking effects.
+(2) We propose a new attack against AIGI detection by exploring the vulnerable frequency region using a post-train Bayesian strategy, avoiding retraining the surrogate model.
+(3) Extensive experiments across 17 AIGI detectors show that our method achieves the highest attack success rates under white-box and black-box settings, outperforming existing baselines.
 II Related Work
-AI-Generated Image Detection: The rapid advancement of AI generative technologies, including GANs [1], VAEs [13], and Diffusion models [2], has brought about substantial risks in the dissemination of disinformation.
-Initially, early VAEs and GANs models were primarily utilized to generate realistic fake face data, leading detection methods [14, 15, 16], to primarily focus on face forgeries.
-However, with the emergence of diffusion models, the scope of generative models has expanded to encompass diverse natural scene objects.
-Consequently, this expansion presents more significant challenges to detection methods, as they must now address a broader range of forgery types and scenarios.
-The primary objective of AIGI detection methods is to discern the authenticity of an image by framing it as a binary classification task, distinguishing between real and fake instances.
-Data-driven-based approaches [3, 17] have demonstrated promising results on generated data identically distributed.
-However, their performance significantly degraded when presented with previously unseen generated images.
-More recently, certain methods [10, 11] focus on analyzing forgery traces in the frequency domain and learning relevant forgery features [18] to enhance the generalization capability of the detection system.
-In pursuit of these objectives, some methods have been proposed to learn local forgery features [19, 20] in the spatial domain.
-AIGI fingerprint representations [21] are acquired through the learning of noise patterns.
-Notably, studies such as [22, 4] have demonstrated the utility of utilizing noise in the frequency domain as a fingerprint for enhancing performance.
-A sequence of detection methodologies [23] has been progressively investigated, relying on the practice of freezing pre-trained models.
-These approaches employ parameter-efficient fine-tuning strategies while integrating additional classifiers.
-For instance, UniFD [24] explores pre-trained vision-language models for AIGI detection, showcasing the effectiveness of leveraging such models in the pursuit of robust forgery detection.
-Furthermore, other researchers have pursued a comprehensive investigation of generalized forgery features from various perspectives, ranging from diffusion reconstruction error [25, 26], adversarial teacher-student discrepancy-aware framework [27] or diffusion noise[28].
-The existing models consistently demonstrate enhancement in detecting AI-generated images.
-However, their behaviors under adversarial attacks have not been thoroughly examined.
-Our work is complementary to existing research by making the attempt to systemically evaluate their robustness to adversarial attacks and suggesting potential improvements.
-Adversarial Attack: Adversarial attacks aim to generate adversarial examples that can fool the target model into predicting wrongly.
-The vulnerability of DNNs has sparked significant concerns in many safety-critical fields, including image classification[29, 30], multimedia communication [31] and human activity recognition [32].
-Very recently, the adversarial attack against Deepfake detection has drawn attention.
-Hussain et al.
-[6] employ gradient sign-based methods [33, 34] to evaluate the adversarial robustness of video Deepfake detectors.
-Li et al.
-[35] generate adversarial fake images along the latent face manifold.
-Carlini et al.
-[5] and Neekhara et al.
-[7] investigate the adversarial threat to Deepfake detection under the black-box setting.
-Jia et al.
-[9] propose to generate adversarial perturbations in the frequency domain.
-Hou et al.
-[8] narrow the feature distribution gap between real and fake images by incorporating natural degradation noise with adversarial perturbations.
-FakePolisher [36] and TraceEvader [37] improve the untraceability of deepfakes by removing artificial fingerprints.
-Different from Deepfake tampering the real face images, AI-generated image encompasses a broader range of synthetic images with more diverse semantic content, amplifying their potential for dissemination of disinformation.
-However, the adversarial robustness of AIGI detection has received little attention.
-In this paper, we propose a universal attack against both AIGI detectors and Deepfake detectors, and evaluate under the white-box and a more practical black-box setting.
+AI-Generated Image Detection: The rapid progress of generative models, from early GANs [1] to recent diffusion models [2], has raised substantial risks of disinformation.
+While early work focused primarily on detecting fake human faces [15, 16, 17, 18], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection.
+AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images.
+Data-driven methods [3] achieve strong in-distribution performance but generalize poorly to unseen generators.
+To address this, frequency-domain approaches [13, 19] exploit forgery traces, while spatial-domain methods learn local forgery features [20].
+Others leverage fingerprint representations from noise patterns [21, 4].
+Parameter-efficient strategies based on frozen pre-trained models have also been explored [22, 23], along with approaches exploiting diffusion reconstruction error [24, 25], teacher–student discrepancies [26], or diffusion noise [27].
+Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [28, 5], multimedia communication [29], and human activity recognition [30].
+Gradient-based attack methods [5, 31] exploit gradients to mislead models.
+To improve adversarial transferability, frequency-domain approaches [32, 6] perturb spectral representations, while ensemble-based attacks [33, 34] aggregate gradients or perturbations from multiple models.
+Very recently, their impact on Deepfake detection has been explored.
+Gradient sign-based attacks [8, 5, 35], latent space manipulations [36], and black-box evaluations [7, 9] reveal vulnerabilities in existing detectors.
+Other works leverage frequency-domain perturbations [11], natural degradation noise [10], or fingerprint removal strategies such as FakePolisher [37] and TraceEvader [38].
+Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [39].
+In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
 III Methodology
@@
 III-B Frequency-based Analysis and Attacks
-Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 21, 4].
-While these subtle clues are invisible in the spatial domain, a series of works [10, 11] demonstrate that there are obvious differences between real and fake images in the frequency domain.
+Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 4].
+While these subtle clues are invisible in the spatial domain, a series of works [13] demonstrate that there are obvious differences between real and fake images in the frequency domain.
 This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective.
@@
 To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [32] to visualize the sensitive components of real and fake images across different models:
-$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
+$\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
 […]
-$\Gamma(\mathbf{x})=\mathcal{D_{I}}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
+$\Gamma(\mathbf{x})=\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
 where $\Gamma(\cdot)$ denotes the random spectrum transformation [32].
@@
 III-D Hybrid Adversarial Attack
-Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3, 21].
+Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3].
 We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains.
@@
 2.
-Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$;$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
+Input: $\mathbf{x}$: training data; $N_{tra}$: the number of training iterations; $M_{\theta^{\prime}}$: sampling iterations for $\theta^{\prime}$$\Theta$: parameters over pre-trained surrogate model; $\{\theta^{\prime}_{1},\dots,\theta^{\prime}_{K}\}$: parameters over appended models; $K$: the number of appended models;
 Output: The adversarial example $\tilde{\mathbf{x}}$;
@@
 $\tilde{\mathbf{x}}^{0}=\mathbf{x}$ ;
-for i = 1 to $I$ do
-       for n = 1 to $N$ do
+for i = 1 to II do
+ for n = 1 to NN do
       Get spectrum transformation output $\Gamma(\tilde{\mathbf{x}}^{i})$ using Eq. 4 ;
@@
 IV-A Experimental Settings
-Datasets: We choose three generated image datasets created by a wide range of generative models, across AIGI detection and Deepfake detection.
+Datasets: We chose three generated image datasets created by a wide range of generative models.
 Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43].
@@
 After training, we collect only the correctly classified testing samples for attack in evaluation.
-Evaluated Models: We extensively evaluate the transferability of adversarial examples by 13 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods.
+TABLE I: Datasets employed for training each detector.
+Synthetic LSUN GenImage Synthetic FFHQ
+CNNSpot
+DenseNet
+EfficientNet
+MobileNet
+Spec
+DCTA
+ViT
+Swin
+GramNet - -
+LGrad - -
+LNP - -
+UnivFD - -
+DNF - -
+FreqNet - -
+FreqMask - -
+DIRE - -
+Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods.
 For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [47], EfficientNet [48] and DenseNet [49], and ViT-based detectors Vision Transformer(ViT) [50] and Swin-Transformer(Swin-ViT) [51].
-For evaluating on various detection methods, we use frequency-based detectors DCTA [11] and Spec [18], gradient-based detectors LGrad [23], CLIP-based detectors UnivFD [24], and diffusion-based detectors DNF [28].
+For evaluating on various detection methods, we use frequency-based detectors DCTA [13], Spec [19], FreqNet [52], and FreqMask [53], gradient-based detectors LGrad [22], CLIP-based detectors UnivFD [23], and diffusion-based detectors DNF [27], DIRE [24] and AEROBLADE [25].
 In addition, LNP [21] extracts the noise pattern of images and GramNet [54] learns the global texture representation.
 We also consider them as victim models.
+The datasets employed for training each detector are summarized in Table I.
+Note that AEROBLADE is a training-free method and not trained on the listed datasets.
 Compared Methods: We adopt gradient-based methods, I-FGSM [5], PGD [31], MI [33] and frequency-based methods S2[32], SSAH [6] and ensemble-based methods ENS [33] and SVRE [34].
-Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks Fakepolisher [36] and TraceEvader [37] as baselines.
+Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [11], Fakepolisher [37] and TraceEvader [38] as baselines.
 Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings.
@@
 Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3].
-Specificaly, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entroy loss with an initial learning rate of 0.0001.
+Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001.
 For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness.
 Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability.
-We transformed images to 224 pixels on LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3].
-On the FFHQ (StyleGAN2) dataset we resized images to 224 pixels to ensure the integrity of the real/fake faces.
+We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3].
+On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces.
 Subsequently, we apply ImageNet normalization across three datasets.
@@
 All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
-TABLE I: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset.
+TABLE II: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic LSUN and GenImage subset.
 “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
@@
 SSAH 0.6 1.0 1.5 99.2 0.5 0.8 1.0 1.1 0.9
-LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
+Synthetic LSUN (ProGAN)
+FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
 IFGSM 78.3 71.5 30.9 34.0 27.1 28.3 11.0 27.8 32.9
@@
 SSAH 1.1 2.0 0.6 96.7 0.5 1.9 1.2 2.2 1.4
-GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
+GenImage(SD)
+FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
 IV-B Attack on Spatial-based and Frequency-based Detectors
-We report the attack performance against spatial-based and frequency-based detectors in Table I.
+We report the attack performance against spatial-based and frequency-based detectors in Table II.
 Under the white-box setting, our proposed method FPBA achieves the highest attack success rate in all cases, and outperforms other competitive methods.
@@
 Note that our proposed FPBA significantly outperforms the SOTA frequency-based attack S2I and SSAH, demonstrating that FPBA is a stronger frequency-based attack against AIGI detectors.
-TABLE II: The attack success rate(%) compared with attacks agaisnt Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.
+TABLE III: The attack success rate(%) compared with attacks against Deepfake detectors(only fake ASR).DenNet, EffNet, MobNet refers to DenseNet, EfficientNet, MobileNet respectively.
 Datasets Attack CNNSpot DenNet EffNet MobNet ViT Swin Ave.
@@
 LSUN TraceEvader 0.6 1.3 4.5 4.8 8.10 11.4 1.9
-(ProGAN) FPBA 100.0 100.0 98.2 99.5 38.6 99.9 89.4
+(ProGAN) FaceAttacker 52.9 30.9 3.7 8.8 2.8 18.9 19.7
+FPBA 100.0 100.0 98.2 99.5 38.6 99.9 89.4
 FakePolisher 93.5 93.2 12.2 95.5 33.5 97.6 70.9
 GenImage TraceEvader 46.1 34.3 9.4 37.0 15.2 44.3 31.1
-(SD) FPBA 100.0 99.7 97.7 98.9 42.1 97.9 89.4
+(SD) FaceAttacker 94.3 79.8 57.9 77.0 12.4 61.2 63.7
+FPBA 100.0 99.7 97.7 98.9 42.1 97.9 89.4
 Figure 5: Visualization of the sensitive frequency components of real and fake images (average 1000 images on LSUN/ProGAN datasets) for frequency-based models.
@@
 IV-B1 Comparison with Attacks against Deepfake Detectors
-TABLE III: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets.
+TABLE IV: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets.
 1 means setting (1), i.e.
@@
 FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
-TABLE IV: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
-Attack GramNet LGrad LNP UnivFD DNF Ave
-MIFGSM 9.7 6.6 38.5 24.2 46.7 25.1
-PGD 67.7 50.2 47.4 11.5 48.3 45.0
-S2I 49.1 49.1 28.3 8.3 43.0 35.6
-FPBA 87.8 50.2 39.5 16.6 49.8 48.9
-TABLE V: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
+TABLE V: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
+Attack GramNet LGrad LNP UnivFD DNF FreqNet FreqMask Ave
+MIFGSM 9.7 6.6 38.5 24.2 46.7 10.5 50.3 26.6
+PGD 67.7 50.2 47.4 11.5 48.3 48.9 77.2 50.2
+S2I 49.1 49.1 28.3 8.3 43.0 45.3 84.2 43.9
+FPBA 87.8 50.2 39.5 16.6 49.8 47.1 98.4 55.6
+TABLE VI: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
 Detector Acc Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
@@
 Attacks against Deepfake detectors are similar to our attacks.
-However, Deepfake mainly manipulating or synthesizing faces, often crafted by GANs.
+However, Deepfake mainly manipulate or synthesize faces, often crafted by GANs.
 In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models.
@@
 These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors.
-For example, techniques like searching for adversarial points on the face manifold [35] or removing facial forgery traces [36, 37] may not be applicable for AIGI detectors.
-To demonstrate this, we compare with FakePolisher [36] and TraceEvader [37], which are both the SOTA attacks for Deepfake detection.
+For example, techniques like searching for adversarial points on the face manifold [36], removing facial forgery traces [37, 38] or simply updating perturbations in the frequency domain to improve attack imperceptibility [11] may not be applicable for AIGI detectors.
+To demonstrate this, we compare with FakePolisher [37], TraceEvader [38] and FaceAttacker [11], which are the SOTA attacks for Deepfake detection.
 Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images.
-As shown in Table II, FPBA significantly outperforms FakePolisher and TraceEvader on both AIGI datasets.
-Although FakePolisher performs better than TraceEvader, but the former largely reduces the image quality, which is very visible and raises suspicion (Fig.
-7).
-We suspect this is because FakePolisher has only learned the face manifold, limiting its ability to reconstruct fine-grained details in AIGI images.
+As shown in Table III, FPBA consistently achieves the highest attack success rates across all detectors and datasets, outperforming prior Deepfake-based attacks by margins ranging from 11.5% to 87.5%.
+FakePolisher performs better than TraceEvader and FaceAttack (albeit still less effective than FPBA), but the former largely reduces the image quality, which is very visible and raises suspicion, as shown in Table XIV and Fig.
+8.
+This is likely because FakePolisher is primarily trained on the face manifold, limiting its ability to reconstruct fine-grained, identity-preserving details in AIGI images.
 IV-B2 Comparison with Ensemble-based Attacks
@@
 Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models.
-We report the results in Table III.
+We report the results in Table IV.
 First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors.
@@
 Nevertheless, FPBA still gets an top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM respectively.
-TABLE VI: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
+TABLE VII: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
 Victim Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
@@
 ViT 19.9 14.5 16.6 3.3 20.1 2.1 12.7
-TABLE VII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
+TABLE VIII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
 Victim ASR Midj. SDv1.4 SDv1.5 ADM Wukong BigGAN Ave
@@
 IV-C1 Attack Various SOTA Detection Methods
-Except for detecting in spatial and frequency domian using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts.
-Specifically, GramNet [53] extracts the global texture representation; UnivFD [24] captures semantic feature using pre-trained CLIP [58]; LNP [22] extracts the noise pattern of images; LGrad [23] extract gradient information through a pre-trained model and DNF [28] leverages diffusion noise features, extracting them through an inverse diffusion process [2].
+Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts.
+Specifically, GramNet [54] extracts the global texture representation; UnivFD [23] captures semantic feature using pre-trained CLIP [56]; LNP [21] extracts the noise pattern of images; LGrad [22] extract gradient information through a pre-trained model; DNF [27] leverages diffusion noise features, extracting them through an inverse diffusion process [2]; FreqNet [52] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [53] introduces a frequency-domain mask strategy for data augmentation during training.
 These significant differences among various detection methods present challenges for the transferability of adversarial examples.
-To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table IV.
-Overall, the transfer success rate is relatively low compared to the results in Table I, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other.
-Nevertheless, FPBA still give the best adversarial transferability.
+To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table V.
+Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other.
+Nevertheless, FPBA still achieves the best adversarial transferability.
+TABLE IX: Adversarial attack on AEROBLADE and DIRE. Average precision (AP) is used to measure the attack performance. CNNSpot is used as the surrogate model.
+[…]
+Clean 80.0 100.0
+MIFGSM 55.9 58.9
+PGD 46.9 43.2
+SSI 54.6 44.5
+FPBA 43.1 40.8
+Very recently, reconstruction error-based methods, such as DIRE [24] and AEROBLADE [25] have emerged as competitive approaches for detecting diffusion-generated images.
+Therefore, we adopt them as victim models in our evaluation.
+As both methods rely on diffusion models, their performance degrades considerably on GAN-generated data.
+Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX.
+We follow [25] to report detection performance using average precision (AP).
+FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
 IV-C2 Attack Cross-Generator Image Detection
@@
 Because Swin-ViT achieves the SOTA results on different subsets [14], we use it as the surrogate model.
-The benign accuracy and attack performance on unseen source data are reported in Table V and Table VI respectively.
+The benign accuracy and attack performance on unseen source data are reported in Table VI and Table VII respectively.
 First, the attack under the white-box setting achieves almost 100% success rate.
@@
 In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them.
-By looking closely at the accuracy/ASR on real images and fake images(Table V,Table VII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack.
+By looking closely at the accuracy/ASR on real images and fake images(Table VI,Table VIII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack.
 Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
-IV-D Attack Performance against Compression in Real Scenarios
-Figure 6: Attack performace on ResNet, Efficient, MobileNet and Swin against JPEG compression on LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
-In real-world scenarios, images are often compressed during propagation, which may also distort adversarial noise patterns.
-Therefore, ensuring the effectiveness of adversarial attacks on compressed images is crucial for simulating real-world scenarios.
+Figure 6: Visualization of the spectrum saliency map (average of adversarial examples generated by FPBA using CNNSpot as the surrogate model) for different AIGI detectors.
+TABLE X: Transfer attack success rate on real samples (REAL ASR) and fake samples(FAKE ASR). The surrogate model is chosen as CNNSpot.
+FreqMask UnivFD GramNet LGrad FreqNet
+REAL ASR 96.7 31.5 76.4 14.6 5.8
+FAKE ASR 100.0 1.7 99.2 85.8 88.4
+ASR 98.4 16.6 87.8 50.2 47.1
+IV-D Adversarial Transferability Analysis
+As shown in Table II and Table V, FPBA exhibit superior adversarial transferability compared to existing attacks.
+To further investigate the underlying reason, we visualize the spectrum saliency map for different AIGI detectors in Fig.
+6.
+The map is computed as the average of adversarial examples generated by FPBA using CNNSpot as the surrogate model.
+Notably, CNNSpot, DenseNet, and Swin ViT share highly similar frequency-sensitive areas, leading to common vulnerabilities in these domains.
+In contrast, ViT exhibits fewer overlapping frequency-sensitive features with CNNSpot, resulting in a comparatively lower transfer success rate.
+Next, considering that the AIGI detection task often suffers from imbalanced recognition accuracy between real and fake samples (as shown in Table VI), we further investigate the attack success rates of FPBA on real images and fake images separately, as shown in Table X.
+Low-level feature detectors, including FreqMask, FreqNet, GramNet, LGrad, mainly rely on frequency, texture, or spectral cues.
+We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples.
+We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [60].
+As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack.
+In contrast, high-level detectors like UnivFD, which rely on global semantic features, exhibit overall lower transferability.
+This suggests that models leveraging generalizable semantic representations are inherently more robust.
+Interestingly, we observe that in UnivFD, fake images are more resistant to adversarial attacks than real ones.
+We will leave the phenomenon to explore in future work.
+TABLE XI: AT-based defenses against PGD attack.
+Dataset Detectors AT-methods Clean Robust
+ProGAN CNNSpot PGD-AT 50.6 49.9
+TRADES 50.1 6.5
+UnivFD PGD-AT 69.6 0
+TRADES 95.5 0
+GenImage CNNSpot PGD-AT 50.7 49.7
+TRADES 55.2 9.4
+UnivFD PGD-AT 49.7 0.7
+TRADES 56.8 0
+IV-E Evaluation against Defense Models
+Evaluation against AT-based Methods
+Next, we investigate the attack performance against defense models.
+As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors.
+In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [31] and TRADES [61].
+However, we found that the training process of AT fails to converge.
+Both AT methods cause a substantial drop in clean accuracy, often to around 50%, which is equivalent to random guessing.
+Robust accuracy against PGD attacks also remains low, in some cases collapsing to nearly 0%.
+This phenomenon can be attributed to the intrinsic data imbalance in AIGI detection: models tend to overfit to low-level synthetic artifacts early in training.
+When AT is applied, these brittle features collapse quickly, and adversarial examples easily cross the decision boundary for both real and fake classes.
+As a result, the model fails to learn discriminative and robust representations, producing near-random predictions on clean data and negligible robustness gains.
+Attack Performance under Compression in Real Scenarios
+Figure 7: Attack performance on ResNet, Efficient, MobileNet and Swin against JPEG compression on Synthetic LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
+Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies.
+Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 14] to utilize Gaussian blurring and JPEG compression as a defense.
+In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios.
+Therefore, ensuring the effectiveness of adversarial attacks on compressed images is also crucial for simulating real-world scenarios.
 To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images.
@@
 The results are reported in Fig.
-6.
+7.
 For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance.
@@
 However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
-TABLE VIII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset.
+Next, we follow the suggestions from [3, 14] to simultaneously utilize Gaussian blurring and JPEG compression as a defense.
+As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
+TABLE XII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset.
 Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively.
@@
 FPBA 100.0 99.2(-0.8)
-IV-D1 Attack Defense Models
-Furthermore, we evaluate FPBA on attacking defense models.
-We first employ adversarial training [38] for evaluation, which is the most popular defense baseline.
-However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection.
-Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 12] to utilize Gaussian blurring and JPEG compression as a defense.
-As shown in Table VIII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
-TABLE IX: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets.
+TABLE XIII: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets.
 “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
@@
 SSAH 1.9 36.4 43.7 98.5 2.0 6.2 0.8 2.4 13.3
-FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
-IV-E Visual Quality Analysis
-TABLE X: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
+FFHQ (Style GAN)
+FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
+IV-F Visual Quality Analysis
+TABLE XIV: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
 […]
@@
 We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig.
-7.
+8.
 Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in.
@@
 Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM.
-As reported in Table X, FPBA outperforms other baselines across all quality assessment by a large margin.
+As reported in Table XIV, FPBA outperforms other baselines across all quality assessment by a large margin.
 This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
-IV-F Additional Performance Analysis
-IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
-In Table I, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [54].
+IV-G Additional Performance Analysis
+IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
+In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [33].
 It indicates the possibility of gradient masking [62].
@@
 Most gradient components of adversarial examples generated by IFGSM (Fig.
-8(a)) and MIFGSM (Fig.
+9(a)) and MIFGSM (Fig.
 9 (c)) tend to stabilize around zero, indicating the vanishing gradient leading to a limited transferability to IFGSM and MIFGSM.
@@
 (a) gradient components of IFGSM; (b) gradient components of PGD; (c) gradient components of MIFGSM; (d) gradient components of MIFGSM with random initialization.
-IV-F2 Evaluation on Deepfake Datasets
+IV-G2 Evaluation on Deepfake Datasets
 As we mention before, there is a big difference between AIGI detection and Deepfake detection.
@@
 To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [1].
-We report the results in Table IX.
+We report the results in Table XIII.
 Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance.
 For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
-TABLE XI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
+TABLE XV: Attack success rate (%) under different frequency transforms on the ProGAN dataset.
+Surrogate Transform CNNSpot DenseNet EfficientNet MobileNet Spec DCTA ViT Swin Avg
+CNNSpot FFT 98.75 98.15 76.05 95.70 19.60 47.40 50.85 94.45 72.62
+Wavelet 99.20 98.10 71.20 94.55 19.30 46.90 50.60 92.65 71.56
+DCT 98.90 98.00 76.40 95.90 19.80 48.00 51.50 94.80 72.91
+MobileNet FFT 31.65 40.00 51.00 99.50 22.35 36.50 12.45 44.80 42.28
+Wavelet 23.04 31.15 32.04 91.40 21.85 36.45 11.20 36.55 35.46
+DCT 32.20 40.40 51.70 99.60 22.60 36.40 12.30 44.60 42.48
+TABLE XVI: Ablation Study on ProGAN dataset. The adversarial samples are crafted from CNNSpot.
 Attack DenseNet MobileNet Spec ViT
@@
 Spatial-frequency 96.2 66.2 47.9 21.0
-TABLE XII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations.
+TABLE XVII: The attack success rates (%) of FPBA on normally trained detectors w.r.t the number $N$ of spectrum transformations.
 “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
-$N$ ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
+NN ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Ave.
 0 64.5 64.0 37.3 60.6 17.5 48.4 40.5 59.7 48.8
@@
 20 99.2 98.7 77.5 96.7 20.7 48.6 51.1 95.6 69.8
-IV-G Ablation Study
+IV-H Ablation Study
+The Selection of Frequency Transforms
+We conduct an ablation study comparing the performance of different frequency transforms, including Fast Fourier Transform (FFT), Wavelet Transform (Wavelet) and Discrete Cosine Transform (DCT).
+The results in Table XV shows that FFT, Walevet and DCT all show high attack performance, with DWT performing the best.
+So we use DCT by default.
 Spatial and Frequency Attack
-We conduct an ablation study in Table XI to investigate the impact of computing gradient in different domains.
+We conduct an ablation study in Table XVI to investigate the impact of computing gradient in different domains.
 In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability.
@@
 In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models.
-The adversarial examples were crafted from ResNet-50 in LSUN(ProGAN) dataset.
+The adversarial examples were crafted from ResNet-50 in Synthetic LSUN(ProGAN) dataset.
 With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper.
-As shown in Table XII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate.
+As shown in Table XVII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate.
 Further, when $N>5$, there is a diminishing gain in attack success rate but with increased computation.
@@
 V Conclusion
-In this paper, we fully investigate the robustness of AIGI detectors against adversarial attacks.
-For the task of AIGI detection, we propose a novel frequency-based post-train Bayesian attack, which explores the vulnerable regions in the frequency domain in a Bayesian manner.
-More broadly, our proposed post-train Bayesian attack strategy extends the Bayesian attack family.
-Extensive experiments have been conducted across models, generators and defense methods under both white-box and black-box settings.
-The following conclusions can be drawn from our investigation.
-  • The current SOTA AIGI detectors are vulnerable to adversarial attacks, while the widely-used defense mechanism adversarial training cannot be directly applied to the AIGI detectors, which brings great challenges to build reliable and robust detectors.
-  • AIGI detectors rely on gradient masking to mitigate the transferability of gradient-based attacks, but the effect of gradient masking can be easily circumvented by a simple random initialization strategy.
-  • The attack success rate is positively associated with the detection accuracy.
-Therefore, evaluating detector robustness on low-accuracy subsets is futile, as there are no effective gradients for attack.
-These findings indicate that building an adversarial robust AIGI detector is still an open problem.
-We hope that our work can be used as a protocol to evaluate the robustness of AIGI detectors and draw more attention to this domain.
-Acknowledgments
-This work was supported in part by the NSF China (No. 62302139) and FRFCU-HFUT (JZ2023HGTA0202, JZ2023HGQA0101).
+In this paper, we investigate the adversarial robustness of AIGI detectors and propose a novel frequency-based post-train Bayesian attack that extends the Bayesian attack family.
+Through extensive experiments across models, generators, and defenses under both white-box and black-box settings, we draw the following conclusions: (1) State-of-the-art AIGI detectors remain highly vulnerable to adversarial attacks.
+Although adversarial training is effective in other domains, it cannot be directly applied in AIGI detection, posing major challenges for building robust detectors.
+(2) Many detectors rely on gradient masking to hinder the transferability of gradient-based attacks, yet the defense is easily bypassed with simple random initialization.
+(3) Attack success correlates positively with detection accuracy, suggesting that robustness evaluation on low-accuracy subsets is uninformative.
+Overall, our findings highlight that constructing adversarially robust AIGI detectors remains an open problem.
+We hope our work serves as both a benchmark and a call to further research in this critical area.

### v5

@@
 First, inspired by the obvious difference between real images and fake images in the frequency domain, we add perturbations under the frequency domain to push the image away from its original frequency distribution.
-Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous AIGI detectors, e.g.
-transferring adversarial examples across CNNs and ViTs.
+Second, we explore the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous AIGI detectors, e.g., transferring adversarial examples across CNNs and ViTs.
 This is achieved by introducing a novel post-train Bayesian strategy that turns a single surrogate into a Bayesian one, capable of simulating diverse victim models using one pre-trained surrogate, without the need for re-training.
 We name our method as Frequency-based Post-train Bayesian Attack, or FPBA.
-Through FPBA, we show that adversarial attack is truly a real threat to AIGI detectors, because FPBA can deliver successful black-box attacks across models, generators, defense methods, and even evade cross-generator detection, which is a crucial real-world detection scenario.
-The code will be shared upon acceptance.
-Index Terms:
-AI-generated Image Detection, Adversarial Examples.
+Through FPBA, we demonstrate that adversarial attacks pose a real threat to AIGI detectors.
+FPBA can deliver successful black-box attacks across various detectors, generators, defense methods, and even evade cross-generator and compressed image detection, which are crucial real-world detection scenarios.
+Our code is available at https://github.com/onotoa/fpba.
 I Introduction
 Figure 1: A high-level illustration of our proposed method.
-The notable progress in generative models, such as GANs [1] and Diffusion models [2], is driving the flourishing development of the image synthesis domain.
+The notable progress in generative models, such as GANs [ffhq] and Diffusion models [adm], is driving the flourishing development of the image synthesis domain.
 These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images.
@@
 Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI).
-Recent state-of-the-art detectors [3, 4] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
+Recent state-of-the-art detectors [wang2020cnn, zhong2023rich] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
 However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real.
 Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging.
-Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [5, 6], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts.
+Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [I-FGSM, luo2022frequency], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts.
 This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints.
 Consequently, transferring adversarial patterns across such diverse representations is far less straightforward.
-Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [7, 8, 9, 10, 11].
+Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [carlini2020evading, hussain2021adversarial, neekhara2021adversarial, hou2023evading, jia2022exploring].
 These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces.
@@
 In this paper, we show that AIGI detectors are vulnerable to adversarial attacks.
-Considering that many works [12, 13] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
+Considering that many works [dzanic2020fourier, frank2020leveraging] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
 As illustrated in Fig.
@@
 Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains.
-Further, Transformer-based detectors have demonstrated outstanding performance in detection AI-generated images[14], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors e.g.
-transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II).
+Further, Transformer-based detectors have demonstrated outstanding performance in detecting AI-generated images[zhu2024genimage], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors, e.g., transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II).
 To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate.
@@
 II Related Work
-AI-Generated Image Detection: The rapid progress of generative models, from early GANs [1] to recent diffusion models [2], has raised substantial risks of disinformation.
-While early work focused primarily on detecting fake human faces [15, 16, 17, 18], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection.
+AI-Generated Image Detection: The rapid progress of generative models, from early GANs [ffhq] to recent diffusion models [adm], has raised substantial risks of disinformation.
+While early work focused primarily on detecting fake human faces [miao2025multi, miao2025mixture], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection.
 AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images.
-Data-driven methods [3] achieve strong in-distribution performance but generalize poorly to unseen generators.
-To address this, frequency-domain approaches [13, 19] exploit forgery traces, while spatial-domain methods learn local forgery features [20].
-Others leverage fingerprint representations from noise patterns [21, 4].
-Parameter-efficient strategies based on frozen pre-trained models have also been explored [22, 23], along with approaches exploiting diffusion reconstruction error [24, 25], teacher–student discrepancies [26], or diffusion noise [27].
-Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [28, 5], multimedia communication [29], and human activity recognition [30].
-Gradient-based attack methods [5, 31] exploit gradients to mislead models.
-To improve adversarial transferability, frequency-domain approaches [32, 6] perturb spectral representations, while ensemble-based attacks [33, 34] aggregate gradients or perturbations from multiple models.
+Data-driven methods [wang2020cnn, feng2025deepfake] achieve strong in-distribution performance but generalize poorly to unseen generators.
+To address this, frequency-domain approaches [miao2022hierarchical, miao2023f] exploit forgery traces, while spatial-domain methods learn local forgery features [guo2023ldfnet].
+Others leverage fingerprint representations from noise patterns [liu2022detecting].
+Parameter-efficient strategies based on frozen pre-trained models have also been explored [tan2023learning], along with approaches exploiting diffusion reconstruction error [ricker2024aeroblade], teacher–student discrepancies [zhu2023gendet], or diffusion noise [zhang2023diffusion].
+Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [ran2024adaptive, I-FGSM], multimedia communication [gao2024deepspoof], computational imaging [liang2025understanding] and human activity recognition [diao2024understanding, diaotasar, diao2021basar].
+Gradient-based attack methods [I-FGSM, pgd] exploit gradients to mislead models.
+To improve adversarial transferability, frequency-domain approaches [luo2022frequency] perturb spectral representations, while ensemble-based attacks [dong2018boosting, svre] aggregate gradients or perturbations from multiple models.
 Very recently, their impact on Deepfake detection has been explored.
-Gradient sign-based attacks [8, 5, 35], latent space manipulations [36], and black-box evaluations [7, 9] reveal vulnerabilities in existing detectors.
-Other works leverage frequency-domain perturbations [11], natural degradation noise [10], or fingerprint removal strategies such as FakePolisher [37] and TraceEvader [38].
-Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [39].
+Gradient sign-based attacks [hussain2021adversarial, I-FGSM, fgsm], latent space manipulations [li2021exploring], and black-box evaluations [carlini2020evading, neekhara2021adversarial] reveal vulnerabilities in existing detectors.
+Other works leverage frequency-domain perturbations [jia2022exploring], natural degradation noise [hou2023evading], or fingerprint removal strategies such as FakePolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader].
+Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [zhou2024stealthdiffusion].
 In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
@@
 We add spatial-frequency adversarial perturbations to AI-generated images in a Bayesian manner, so that they are misclassified as real.
-DCT and IDCT are discrete cosine transformation and inverse discrete cosine transformation respectively.
+DCT and IDCT are the discrete cosine transformation and inverse discrete cosine transformation, respectively.
 III-A Preliminaries
@@
 We aim to inject adversarial perturbation into the original image that makes the detector misclassify.
-Such an adversary problem can be optimized by minimizing the predictive probability, i.e.
-maximizing the classification loss:
+Such an adversary problem can be optimized by minimizing the predictive probability, i.e., maximizing the classification loss:
 $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\tilde{\mathbf{x}},\Theta)=\operatorname*{arg\,max}_{\tilde{\mathbf{x}}}L(\tilde{\mathbf{x}},y,\Theta),\text{ s.t.
 }\left\|\delta\right\|_{p}\leq\epsilon,$ (1)
-where $L$ is the binary cross entropy loss in AI-generated image detection.
+where $L$ is the binary cross-entropy loss in AI-generated image detection.
 Adversarial example $\tilde{\mathbf{x}}=\mathbf{x}+\delta$, in which $\delta$ is the adversarial perturbation and $\epsilon$ is the perturbation budget.
 Eq.
-1 can be performed with iterative gradient-based methods, such as PGD [31] or I-FGSM [5]:
+1 can be performed with iterative gradient-based methods, such as PGD [pgd] or I-FGSM [I-FGSM]:
 $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
@@
 III-B Frequency-based Analysis and Attacks
-Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [3, 4].
-While these subtle clues are invisible in the spatial domain, a series of works [13] demonstrate that there are obvious differences between real and fake images in the frequency domain.
+Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [wang2020cnn, zhong2023rich].
+While these subtle clues are invisible in the spatial domain, a series of works [frank2020leveraging] demonstrate that there are obvious differences between real and fake images in the frequency domain.
 This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective.
-To this end, we first implement discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain.
-To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [32] to visualize the sensitive components of real and fake images across different models:
+To this end, we first implement the discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain.
+To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [long2022frequency] to visualize the sensitive components of real and fake images across different models:
 $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
@@
 $\Gamma(\mathbf{x})=\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
-where $\Gamma(\cdot)$ denotes the random spectrum transformation [32].
+where $\Gamma(\cdot)$ denotes the random spectrum transformation [long2022frequency].
 $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$.
@@
 3(a), tuning the spectrum saliency map can cover most of the other models.
-We hence conduct adversarial attack in the frequency domain via spectrum transformation:
+We hence conduct adversarial attacks in the frequency domain via spectrum transformation:
 $\operatorname*{arg\,min}_{\tilde{\mathbf{x}}}p(y\mid\Gamma(\tilde{\mathbf{x}}),\Theta),\text{ s.t. }\left\|\delta\right\|_{p}\leq\epsilon.$ (5)
@@
 Although tuning the spectrum transformation in Eq.
-5 can simulate different substitute models with a homogeneous architecture [32], it shows limited transferability when applied to heterogeneous architectures, e.g.
-transferring adversarial examples across ViTs and CNNs.
-This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e.
-exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models.
+5 can simulate different substitute models with a homogeneous architecture [long2022frequency], it shows limited transferability when applied to heterogeneous architectures, e.g., transferring adversarial examples across ViTs and CNNs.
+This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e., exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models.
 Therefore, we redefine Eq.
@@
 However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors.
-First, the Bayesian posterior for DNNs is a high dimensional distribution due to a very large number of parameters of DNNs [40].
+First, the Bayesian posterior for DNNs is a high-dimensional distribution due to a very large number of parameters of DNNs [izmailov2021bayesian].
 Hence computing and sampling the posterior distribution is an intractable problem.
-Albeit feasible for approximate sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high dimensional space.
-Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [14, 41].
+Albeit feasible for approximately sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high-dimensional space.
+Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [zhu2024genimage, he2021forgerynet].
 From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
@@
 III-C2 Inference on Bayesian Appended Models
-$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[42] to sample appended model $\theta^{\prime}$ in each iteration:
+$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[springenberg_bayesian_2016] to sample appended model $\theta^{\prime}$ in each iteration:
 $\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I}),$
 $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2},$ (10)
-where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [42].
+where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [springenberg_bayesian_2016].
 III-D Hybrid Adversarial Attack
-Despite detecting fingerprints fake images in the frequency domain, some works also extract fingerprint features in the spatial domain [3].
+Despite detecting fake fingerprints in the frequency domain, some works also extract fingerprint features in the spatial domain [wang2020cnn].
 We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains.
@@
 Datasets: We chose three generated image datasets created by a wide range of generative models.
-Synthetic LSUN is a commonly used dataset proposed by CNNSpot [3], containing 360k real images from LSUN and 360k fake images generated by ProGAN [43].
-GenImage [14] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
-Following the protocol in [14], we employ a subset of GenImage collecting 162k real images from Imagenet [44] and 162k Stable Diffusion(SD) V1.4 [45] generated images for training.
+Synthetic LSUN is a commonly used dataset proposed by CNNSpot [wang2020cnn], containing 360k real images from LSUN and 360k fake images generated by ProGAN [progan].
+GenImage [zhu2024genimage] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
+Following the protocol in [zhu2024genimage], we employ a subset of GenImage, collecting 162k real images from Imagenet [deng2009imagenet] and 162k Stable Diffusion(SD) V1.4 [sd] generated images for training.
 The images generated by the other generators are used for testing in Section IV-C2.
-To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [46].
-The Deepfake dataset consists of 50k real face images from FFHQ [1] and 50k generated face images generated by StyleGAN2 [1].
+To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [shamshad2023evading].
+The Deepfake dataset consists of 50k real face images from FFHQ [ffhq] and 50k generated face images generated by StyleGAN2 [ffhq].
 After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
 Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods.
-For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [3], MobileNet [47], EfficientNet [48] and DenseNet [49], and ViT-based detectors Vision Transformer(ViT) [50] and Swin-Transformer(Swin-ViT) [51].
-For evaluating on various detection methods, we use frequency-based detectors DCTA [13], Spec [19], FreqNet [52], and FreqMask [53], gradient-based detectors LGrad [22], CLIP-based detectors UnivFD [23], and diffusion-based detectors DNF [27], DIRE [24] and AEROBLADE [25].
-In addition, LNP [21] extracts the noise pattern of images and GramNet [54] learns the global texture representation.
+For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [wang2020cnn], MobileNet [howard2017mobilenets], EfficientNet [tan2019efficientnet] and DenseNet [huang2017densely], and ViT-based detectors Vision Transformer(ViT) [vit] and Swin-Transformer(Swin-ViT) [liu2021swin].
+For evaluating on various detection methods, we use frequency-based detectors DCTA [frank2020leveraging], Spec [zhang2019detecting], FreqNet [tan2024frequency], and FreqMask [doloriel2024frequency], gradient-based detectors LGrad [tan2023learning], CLIP-based detectors UnivFD [ojha2023towards], and diffusion-based detectors DNF [zhang2023diffusion], DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade].
+In addition, LNP [liu2022detecting] extracts the noise pattern of images and GramNet [GramNet] learns the global texture representation.
 We also consider them as victim models.
@@
 Note that AEROBLADE is a training-free method and not trained on the listed datasets.
-Compared Methods: We adopt gradient-based methods, I-FGSM [5], PGD [31], MI [33] and frequency-based methods S2[32], SSAH [6] and ensemble-based methods ENS [33] and SVRE [34].
-Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [11], Fakepolisher [37] and TraceEvader [38] as baselines.
+Compared Methods: We adopt gradient-based methods, I-FGSM [I-FGSM], PGD [pgd], MI [dong2018boosting] and frequency-based methods S2[long2022frequency], SSAH [luo2022frequency] and ensemble-based methods ENS [dong2018boosting] and SVRE [svre].
+Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [jia2022exploring], Fakepolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader] as baselines.
 Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings.
 For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
-Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [3].
+Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [wang2020cnn].
 Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001.
-For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [3] to improve the models’ generalization and robustness.
+For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [wang2020cnn] to improve the models’ generalization and robustness.
 Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability.
-We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [3].
+We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [wang2020cnn].
 On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces.
@@
 For other AIGI detectors, we use the pre-trained model from their official code.
-For post-train Bayesian optimization, we follow the default setting in [55].
+For post-train Bayesian optimization, we follow the default setting in [bbc].
 Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate.
 Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$.
-For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [32].
+For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [long2022frequency].
 All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
 SSAH 0.6 1.0 1.5 99.2 0.5 0.8 1.0 1.1 0.9
-Synthetic LSUN (ProGAN)
-FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
+Synthetic LSUN (ProGAN) FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
 IFGSM 78.3 71.5 30.9 34.0 27.1 28.3 11.0 27.8 32.9
@@
 SSAH 1.1 2.0 0.6 96.7 0.5 1.9 1.2 2.2 1.4
-GenImage(SD)
-FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
+GenImage(SD) FPBA(Ours) 49.7 46.8 44.0 100 44.1 49.4 16.6 45.0 42.2
 IV-B Attack on Spatial-based and Frequency-based Detectors
@@
 TABLE IV: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets.
-1 means setting (1), i.e.
-taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e.
-taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e.
-taking an ensemble of CNNSpot, DCTA and ViT.
-4 means only using CNNSpot as surrogate model.
+1 means setting (1), i.e., taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e., taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e., taking an ensemble of CNNSpot, DCTA and ViT.
+4 means only using CNNSpot as a surrogate model.
 Attack CNNSpot DenNet EffNet MobNet Spec DCTA ViT Swin Ave
@@
 FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
-TABLE V: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
+TABLE V: Transfer-based attack against SOTA detectors on Synthetic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [wang2020cnn].
 Attack GramNet LGrad LNP UnivFD DNF FreqNet FreqMask Ave
@@
 Attacks against Deepfake detectors are similar to our attacks.
-However, Deepfake mainly manipulate or synthesize faces, often crafted by GANs.
+However, Deepfakes mainly manipulate or synthesize faces, often crafted by GANs.
 In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models.
@@
 These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors.
-For example, techniques like searching for adversarial points on the face manifold [36], removing facial forgery traces [37, 38] or simply updating perturbations in the frequency domain to improve attack imperceptibility [11] may not be applicable for AIGI detectors.
-To demonstrate this, we compare with FakePolisher [37], TraceEvader [38] and FaceAttacker [11], which are the SOTA attacks for Deepfake detection.
+For example, techniques like searching for adversarial points on the face manifold [li2021exploring], removing facial forgery traces [huang2020fakepolisher, wu2024traceevader] or simply updating perturbations in the frequency domain to improve attack imperceptibility [jia2022exploring] may not be applicable for AIGI detectors.
+To demonstrate this, we compare with FakePolisher [huang2020fakepolisher], TraceEvader [wu2024traceevader] and FaceAttacker [jia2022exploring], which are the SOTA attacks for Deepfake detection.
 Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images.
@@
 IV-B2 Comparison with Ensemble-based Attacks
-Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [33] and SVRE [34], i.e.
-utilizing an ensemble of surrogate models to generate adversarial examples.
+Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [dong2018boosting] and SVRE [svre], i.e., utilizing an ensemble of surrogate models to generate adversarial examples.
 To examine the impact of different surrogate combinations(CNN-based vs.
@@
 We report the results in Table IV.
-First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors.
+First, compared with setting (1), the heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors.
 We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate.
@@
 Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases.
-For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet.
+For instance, the adversarial examples only generated by CNNSpot with FPBA get a success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet.
 This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
@@
 Therefore, the gradient information available from the spatial domain is much richer compared to the frequency domain.
-Nevertheless, FPBA still gets an top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM respectively.
+Nevertheless, FPBA still gets a top-2 average transfer success rate of 43.1% against frequency-based detectors, which is only slightly lower 1.5% than PGD, and much higher 3.3%, 15.8% and 17.6% than S2I, MIFGSM and IFGSM, respectively.
 TABLE VII: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
@@
 Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts.
-Specifically, GramNet [54] extracts the global texture representation; UnivFD [23] captures semantic feature using pre-trained CLIP [56]; LNP [21] extracts the noise pattern of images; LGrad [22] extract gradient information through a pre-trained model; DNF [27] leverages diffusion noise features, extracting them through an inverse diffusion process [2]; FreqNet [52] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [53] introduces a frequency-domain mask strategy for data augmentation during training.
+Specifically, GramNet [GramNet] extracts the global texture representation; UnivFD [ojha2023towards] captures semantic feature using pre-trained CLIP [CLIP]; LNP [liu2022detecting] extracts the noise pattern of images; LGrad [tan2023learning] extract gradient information through a pre-trained model; DNF [zhang2023diffusion] leverages diffusion noise features, extracting them through an inverse diffusion process [adm]; FreqNet [tan2024frequency] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [doloriel2024frequency] introduces a frequency-domain mask strategy for data augmentation during training.
 These significant differences among various detection methods present challenges for the transferability of adversarial examples.
 To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table V.
-Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other.
+Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability between each other.
 Nevertheless, FPBA still achieves the best adversarial transferability.
@@
 FPBA 43.1 40.8
-Very recently, reconstruction error-based methods, such as DIRE [24] and AEROBLADE [25] have emerged as competitive approaches for detecting diffusion-generated images.
+Very recently, reconstruction error-based methods, such as DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade] have emerged as competitive approaches for detecting diffusion-generated images.
 Therefore, we adopt them as victim models in our evaluation.
@@
 Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX.
-We follow [25] to report detection performance using average precision (AP).
+We follow [ricker2024aeroblade] to report detection performance using average precision (AP).
 FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
 IV-C2 Attack Cross-Generator Image Detection
-One important real-world detection problem is cross-generator image detection, i.e.
-identifying fake images generated by unseen generative models.
+One important real-world detection problem is cross-generator image detection, i.e., identifying fake images generated by unseen generative models.
 We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection.
-We train the detectors on images generated by SD v1.4 [45] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [57], SD V1.4 [45], SD V1.5 [45], ADM [2], Wukong [58] and BigGAN [59].
-Because Swin-ViT achieves the SOTA results on different subsets [14], we use it as the surrogate model.
+We train the detectors on images generated by SD v1.4 [sd] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [Midjourney], SD V1.4 [sd], SD V1.5 [sd], ADM [adm], Wukong [wukong] and BigGAN [biggan].
+Because Swin-ViT achieves the SOTA results on different subsets [zhu2024genimage], we use it as the surrogate model.
 The benign accuracy and attack performance on unseen source data are reported in Table VI and Table VII respectively.
-First, the attack under the white-box setting achieves almost 100% success rate.
-Second, the transfer success rate is positively correlated to the accuracy on unseen source data.
+First, the attack under the white-box setting achieves an almost 100% success rate.
+Second, the transfer success rate is positively correlated with the accuracy on unseen source data.
 The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high.
@@
 We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples.
-We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [60].
+We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [yanorthogonal].
 As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack.
@@
 As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors.
-In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [31] and TRADES [61].
+In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [pgd] and TRADES [trades].
 However, we found that the training process of AT fails to converge.
@@
 Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies.
-Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 14] to utilize Gaussian blurring and JPEG compression as a defense.
+Considering that AIGI detection typically [wang2020cnn] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [wang2020cnn, zhu2024genimage] to utilize Gaussian blurring and JPEG compression as a defense.
 In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios.
@@
 For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases.
-However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
-Next, we follow the suggestions from [3, 14] to simultaneously utilize Gaussian blurring and JPEG compression as a defense.
+However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under real-world scenarios.
+Next, we follow the suggestions from [wang2020cnn, zhu2024genimage] to simultaneously utilize Gaussian blurring and JPEG compression as a defense.
 As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@
 SSAH 1.9 36.4 43.7 98.5 2.0 6.2 0.8 2.4 13.3
-FFHQ (Style GAN)
-FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
+FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
 IV-F Visual Quality Analysis
@@
 Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in.
-In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images.
-In contrast, FPBA generate more natural-looking adversarial examples.
-Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM.
-As reported in Table XIV, FPBA outperforms other baselines across all quality assessment by a large margin.
-This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
+In addition, the fine-grained details of objects generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images.
+In contrast, FPBA generates more natural-looking adversarial examples.
+Further, we also report the quantitative results using common metrics for image quality assessment, including MSE, PSNR(db) and SSIM.
+As reported in Table XIV, FPBA outperforms other baselines across all quality assessments by a large margin.
+This suggests that FPBA adding adversarial noise in the frequency domain, rather than directly in the spatial domain, is more imperceptible to observers.
 IV-G Additional Performance Analysis
 IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
-In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [33].
-It indicates the possibility of gradient masking [62].
+In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have a higher success rate [dong2018boosting].
+It indicates the possibility of gradient masking [athalye2018obfuscated].
 To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties.
@@
 IV-G2 Evaluation on Deepfake Datasets
-As we mention before, there is a big difference between AIGI detection and Deepfake detection.
+As we mentioned before, there is a big difference between AIGI detection and Deepfake detection.
 But we empirically found that FPBA is also effective for face forgery detection.
-To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [46], which is consist of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [1].
+To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [shamshad2023evading], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [ffhq].
 We report the results in Table XIII.
-Although FPBA is not specific designed for Deepfake detection, it still achieve the best white-box attack performance.
+Although FPBA is not specifically designed for Deepfake detection, it still achieves the best white-box attack performance.
 For transfer-based attack, FPBA also has the top-2 performance, demonstrating that FPBA is a universal threat across both AIGI and Deepfake detection.
@@
 We conduct an ablation study comparing the performance of different frequency transforms, including Fast Fourier Transform (FFT), Wavelet Transform (Wavelet) and Discrete Cosine Transform (DCT).
-The results in Table XV shows that FFT, Walevet and DCT all show high attack performance, with DWT performing the best.
+The results in Table XV show that FFT, Walevet and DCT all show high attack performance, with DWT performing the best.
 So we use DCT by default.
 Spatial and Frequency Attack
-We conduct an ablation study in Table XVI to investigate the impact of computing gradient in different domains.
-In comparison with computing attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability.
+We conduct an ablation study in Table XVI to investigate the impact of computing the gradient in different domains.
+In comparison with computing the attack gradient solely in the spatial domain or frequency domain, our spatial-frequency attack achieves higher transferability.
 More ablation studies can be found in the supplementary material.
@@
 In this study, we investigate the impacts on the number ($N$) of spectrum transformations, which can simulate diverse substitute models.
-The adversarial examples were crafted from ResNet-50 in Synthetic LSUN(ProGAN) dataset.
-With the exception of $N$, other hyper-parameters keep the same with the default settings in the paper.
+The adversarial examples were crafted from ResNet-50 in the Synthetic LSUN(ProGAN) dataset.
+With the exception of $N$, other hyperparameters keep the same with the default settings in the paper.
 As shown in Table XVII, with $N$ increasing 5 from 0, FPBA shows a substantial augmentation in attack success rate.
@@
 We hope our work serves as both a benchmark and a call to further research in this critical area.
+Acknowledgments
+This work was supported in part by the NSF China (No. 62302139, 62576020) and Fundamental Research Funds for the Central Universities of China (PA2025IISL0113, JZ2025HGTB0227).
+[Uncaptioned image] Yunfeng Diao is a Lecturer in the School of Computer Science and Information Engineering, Hefei University of Technology, China.
+He received his PhD from Southwest Jiaotong University, China.
+His current research interests include computer vision and the security of machine learning.
+He has published over 30 papers in leading venues such as CVPR, ICLR, ICCV, ICML, IEEE TMM and IEEE TCSVT.
+He has been an editorial board member of IJAACS, the chief organizer of several workshops at IJCAI, also regularly reviewing for top-tier journals and conferences.
+He has received the Outstanding Contribution Award at the IJCAI Workshop and the Best Paper Award at the IROS RODGE Workshop.
+[Uncaptioned image] Naixin Zhai received the B.S.
+degree from Hefei University of Technology, Hefei, China, in 2024.
+He is current working toward the M.S.
+degree with the Department of Electronic Engineering and Information Science.
+His research interests include deep learning and synthetic image detection.
+[Uncaptioned image] Changtao Miao received his B.S.
+degree in 2019 from AnHui University.
+He is currently pursuing the Ph.D.
+degree in Cyber Science and Technology in University of Science and Technology of China.
+His research interests include face forgery forensics and face manipulation.
+[Uncaptioned image] Zitong Yu (Senior Member, IEEE) received the Ph.D.
+degree in computer science and engineering from the University of Oulu, Finland, in 2022.
+Currently, he is an Assistant Professor with Great Bay University, China.
+He was a Post-Doctoral Researcher with the ROSE Laboratory, Nanyang Technological University.
+He was a Visiting Scholar with TVG, University of Oxford, from July 2021 to November 2021.
+His research interests include human-centric computer vision and biometric security.
+He was a recipient of IAPR Best Student Paper Award, IEEE Finland Section Best Student Conference Paper Award 2020.
+[Uncaptioned image] Xingxing Wei received his Ph.D degree in computer science from Tianjin University, and B.S.
+degree in Automation from Beihang University (BUAA), China.
+He is now an Associate Professor at Beihang University (BUAA).
+His research interests include computer vision, adversarial machine learning and its applications to multimedia content analysis.
+He is the author of referred journals and conferences in IEEE TPAMI, TMM, TCYB, TGRS, IJCV, PR, CVIU, CVPR, ICCV, ECCV, ACMMM, AAAI, IJCAI etc.
+[Uncaptioned image] Xun Yang (Senior Member, IEEE) received the Ph.D.
+degree from the Hefei University of Technology, Hefei, China, in 2017.
+He is currently a Professor with the Department of Electronic Engineering and Information Science, University of Science and Technology of China (USTC).
+From 2015 to 2017, he visited the University of Technology Sydney (UTS), Australia, as a Joint Ph.D.
+Student.
+He was a Research Fellow with the NExT++ Research Center, National University of Singapore (NUS), from 2018 to 2021.
+His current research interests include information retrieval, cross-media analysis and reasoning, and computer vision.
+He serves as an Associate Editor for IEEE TRANSACTIONS ON BIG DATA and Multimedia Systems journal.
+[Uncaptioned image] Meng Wang (Fellow, IEEE) is a professor at Hefei University of Technology, China.
+His current research interests include multimedia content analysis, computer vision, and pattern recognition.
+He received paper prizes or awards from ACM MM 2009 (Best Paper Award), ACM MM 2010 (Best Paper Award), MMM 2010 (Best Paper Award), SIGIR 2015 (Best Paper Honorable Mention), IEEE TMM 2015 and 2016 (Prize Paper Award Honorable Mention) and ACM TOMM 2018 (Nicolas D.
+Georganas Best Paper Award), etc.
+He currently serves on the editorial/advisory boards of IEEE TPAMI, IEEE TMM, IEEE TNNLS, etc.
+He is a fellow of IEEE and IAPR.
+Vulnerabilities in AI-generated Image Detection:
+The Challenge of Adversarial Attacks
+—Supplementary Document—
+Figure 10: Visualization of the spectrum saliency map with different models.
+VI Additional Experiments
+VI-A Experiment Details
+We compared a series of distinguished backbone architectures pretrained on the ImageNet dataset.
+Specifically, we evaluated ResNet-50, DenseNet-121, EfficientNet-B4, MobileNet-V2, ViT-B/16, and Swin-B.
+Additionally, we examined Spec equipped with ResNet-34 architecture, as well as DCTA equipped with ResNet-50 architecture.
+In Table XVIII, we enumerate the accuracy of all models on the test set as observed in our experiments.
+The meticulous evaluation of these foundational models not only revealed the potential vulnerabilities of AIGI detection models but also highlighted the adversarial threats they face.
+Our analytical results offer significant insights for understanding and improving the robustness of AIGI detection models.
+TABLE XVIII: The accuracy of each detector on different datasets.
+Model ResNet DenNet EffNet MobNet Spec DCTA ViT Swin
+LSUN(ProGAN) 99.99 99.91 99.96 99.93 99.90 99.91 99.74 100.00
+GenImage(SD) 99.55 99.92 99.9 99.93 99.12 99.17 99.48 99.97
+FFHQ(Style GAN) 99.81 99.89 99.83 99.63 96.51 96.6 99.55 99.99
+VI-B Evaluation on ResNeXt
+We use ResNext [xie2017aggregated] as the surrogate to evaluate the attack performance.
+As reported in Table XIX, FPBA still achieves the best average attack success rate, which is consistent with other surrogate models.
+TABLE XIX: Attack evaluation using ResNeXt as surrogate model on the ProGAN-LSUN dataset.
+Attack ResNeXt ResNet DenNet Spec DCTA ViT Swin Ave.
+MI-FGSM 54.2 29.7 21.3 24.8 49.7 14.2 6.6 28.6
+PGD 77.1 35.4 30.9 28.0 50.4 20.6 28.0 38.6
+S2I 98.6 90.6 86.9 13.4 36.9 15.0 68.0 58.5
+FPBA 99.9 89.0 89.2 21.5 49.0 29.6 64.1 63.2
+VI-C Study of Attack Strength
+As shown in Table XX, we report the comparison of the attack success rates of our FPBA with existing methods under different attack strengths.
+The results reflect a pattern consistent with when $\epsilon$=8/255, where our method can achieve globally optimal results compared to previous methods, and a larger perturbation budget consistently enhances adversarial transferability across various surrogate models.
+However, as the attack strength increases, the noise added causes more and more damage to the image, and the visual impact becomes greater, which is revealed in Fig.
+11.
+VI-D Spectrum Saliency Map
+In Fig. 10, we present the spectrum saliency map of all detectors for the original samples and the spectrum saliency map after frequency domain transformation.
+TABLE $X$: The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN(ProGAN) with $\epsilon$=4/255, 8/255 or 12/255.
+“Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
+Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
+IFGSM 52.1 44.6 14.0 36.8 15.0 41.2 12.9 24.6 27.0
+MIFGSM 52.1 50.0 23.5 45.4 15.0 45.1 22.9 38.3 34.3
+ResNet-50 PGD 53.5 46.3 15.5 38.7 20.3 44.4 12.4 26.0 29.1
+S2I 97.1 68.1 21.1 33.0 7.0 18.3 2.0 17.9 23.0
+$\epsilon$=4/255 FPBA(Ours) 98.9 93.7 55.0 79.8 15.8 40.8 18.6 75.1 54.1
+IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
+MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
+ResNet-50 PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
+S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
+$\epsilon$=8/255 FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
+IFGSM 52.1 51.9 28.7 50.0 18.3 49.1 40.6 50.2 41.2
+MIFGSM 52.1 51.8 28.7 47.4 37.2 50.0 34.7 38.9 41.2
+ResNet-50 PGD 96.3 96.1 67.0 91.9 30.6 48.8 58.6 87.5 68.6
+S2I 97.3 88.4 78.5 89.3 28.5 47.6 25.4 85.3 63.3
+$\epsilon$=12/255 FPBA(Ours) 98.9 99.3 82.9 98.5 20.7 49.2 68.7 98.0 73.9
+IFGSM 4.0 5.6 12.5 75.4 12.3 24.6 4.8 3.4 9.6
+MIFGSM 9.0 11.5 16.5 75.4 13.7 26.4 6.4 10.9 13.5
+MobileNet-V2 PGD 3.4 5.0 11.8 87.7 15.1 25.7 4.5 3.2 9.8
+S2I 6.0 7.0 19.2 93.6 5.9 8.0 1.2 5.6 7.6
+$\epsilon$=4/255 FPBA(Ours) 7.9 11.8 26.0 99.5 14.4 25.8 4.5 9.5 14.3
+IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
+MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
+MobileNet-V2 PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
+S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
+$\epsilon$=8/255 FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
+IFGSM 18.8 21.4 27.7 75.4 18.0 40.2 11.4 23.0 22.9
+MIFGSM 25.2 25.2 33.7 75.4 22.5 48.7 21.6 25.2 28.9
+MobileNet-V2 PGD 34.8 39.8 27.6 99.8 35.4 48.5 14.7 45.1 35.1
+S2I 28.1 33.5 44.0 98.6 23.3 34.5 7.9 35.5 29.5
+$\epsilon$=12/255 FPBA(Ours) 42.3 45.7 56.3 99.9 27.0 42.0 18.7 49.0 40.1
+Figure 11: The visual results of adversarial examples with different perturbation budgets.

### v6

@@
 Figure 1: A high-level illustration of our proposed method.
-The notable progress in generative models, such as GANs [ffhq] and Diffusion models [adm], is driving the flourishing development of the image synthesis domain.
+The notable progress in generative models, such as GANs [27] and Diffusion models [5], is driving the flourishing development of the image synthesis domain.
 These generated fake images exhibit realistic-looking, rendering them visually indistinguishable from real images.
@@
 Consequently, numerous detectors have been proposed to identify AI-generated images (AIGI).
-Recent state-of-the-art detectors [wang2020cnn, zhong2023rich] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
+Recent state-of-the-art detectors [54, 63] rely on Deep Neural Networks (DNNs) to classify, achieving significant accuracy performance across multiple datasets and generative models.
 However, our investigation reveals that AIGI detectors are vulnerable to adversarial examples, capable of misleading detectors by classifying fake images as real.
 Albeit identifying a key issue that needs to be addressed, designing an effective attack for AIGI detection is still challenging.
-Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [I-FGSM, luo2022frequency], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts.
+Unlike image classification, which primarily investigates adversarial vulnerability in high-level semantic representations shared across models [28, 35], AIGI detection relies on discriminative information from both high-level semantics and subtle low-level generative artifacts.
 This combination produces a more diverse and heterogeneous feature space, where adversarial perturbations must simultaneously disrupt semantic cues and model-specific generative fingerprints.
 Consequently, transferring adversarial patterns across such diverse representations is far less straightforward.
-Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [carlini2020evading, hussain2021adversarial, neekhara2021adversarial, hou2023evading, jia2022exploring].
+Next, several works have explored the adversarial robustness in GAN-based Deepfake detection [3, 23, 42, 19, 25].
 These methods largely focus on identifying facial fingerprints unique to GANs and exploiting real-fake differences in manipulated human faces.
@@
 In this paper, we show that AIGI detectors are vulnerable to adversarial attacks.
-Considering that many works [dzanic2020fourier, frank2020leveraging] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
+Considering that many works [12, 14] have demonstrated the obvious changes between real and fake images in the frequency domain, we explore the vulnerable region of AIGI detectors in the frequency domain.
 As illustrated in Fig.
@@
 Therefore, we utilize frequency spectrum transformation to uncover diverse substitute models via adding adversarial perturbations in the various frequency transformation domains.
-Further, Transformer-based detectors have demonstrated outstanding performance in detecting AI-generated images[zhu2024genimage], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors, e.g., transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II).
+Further, Transformer-based detectors have demonstrated outstanding performance in detecting AI-generated images[66], but we have observed that there is an obvious gap in adversarial transferability across heterogeneous AIGI detectors, e.g., transferring adversarial examples across Convolutional Neural Networks (CNNs) to Visual Transformers (ViTs) (as shown in Table II).
 To tackle this issue, we propose a post-train Bayesian strategy to conduct a Bayesian treatment on the surrogate model, without the need for re-training the surrogate.
@@
 II Related Work
-AI-Generated Image Detection: The rapid progress of generative models, from early GANs [ffhq] to recent diffusion models [adm], has raised substantial risks of disinformation.
-While early work focused primarily on detecting fake human faces [miao2025multi, miao2025mixture], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection.
+AI-Generated Image Detection: The rapid progress of generative models, from early GANs [27] to recent diffusion models [5], has raised substantial risks of disinformation.
+While early work focused primarily on detecting fake human faces [38, 37], the versatility of diffusion models has broadened generation to diverse scenes, posing greater challenges for detection.
 AIGI detection is typically formulated as a binary classification task to distinguish real from synthetic images.
-Data-driven methods [wang2020cnn, feng2025deepfake] achieve strong in-distribution performance but generalize poorly to unseen generators.
-To address this, frequency-domain approaches [miao2022hierarchical, miao2023f] exploit forgery traces, while spatial-domain methods learn local forgery features [guo2023ldfnet].
-Others leverage fingerprint representations from noise patterns [liu2022detecting].
-Parameter-efficient strategies based on frozen pre-trained models have also been explored [tan2023learning], along with approaches exploiting diffusion reconstruction error [ricker2024aeroblade], teacher–student discrepancies [zhu2023gendet], or diffusion noise [zhang2023diffusion].
-Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [ran2024adaptive, I-FGSM], multimedia communication [gao2024deepspoof], computational imaging [liang2025understanding] and human activity recognition [diao2024understanding, diaotasar, diao2021basar].
-Gradient-based attack methods [I-FGSM, pgd] exploit gradients to mislead models.
-To improve adversarial transferability, frequency-domain approaches [luo2022frequency] perturb spectral representations, while ensemble-based attacks [dong2018boosting, svre] aggregate gradients or perturbations from multiple models.
+Data-driven methods [54, 13] achieve strong in-distribution performance but generalize poorly to unseen generators.
+To address this, frequency-domain approaches [40, 39] exploit forgery traces, while spatial-domain methods learn local forgery features [17].
+Others leverage fingerprint representations from noise patterns [31].
+Parameter-efficient strategies based on frozen pre-trained models have also been explored [51], along with approaches exploiting diffusion reconstruction error [46], teacher–student discrepancies [65], or diffusion noise [62].
+Adversarial Attack: Adversarial attacks craft perturbed examples that mislead target models, raising concerns in safety-critical applications such as image classification [45, 28], multimedia communication [15], computational imaging [30] and human activity recognition [7, 8, 6].
+Gradient-based attack methods [28, 36] exploit gradients to mislead models.
+To improve adversarial transferability, frequency-domain approaches [35] perturb spectral representations, while ensemble-based attacks [10, 58] aggregate gradients or perturbations from multiple models.
 Very recently, their impact on Deepfake detection has been explored.
-Gradient sign-based attacks [hussain2021adversarial, I-FGSM, fgsm], latent space manipulations [li2021exploring], and black-box evaluations [carlini2020evading, neekhara2021adversarial] reveal vulnerabilities in existing detectors.
-Other works leverage frequency-domain perturbations [jia2022exploring], natural degradation noise [hou2023evading], or fingerprint removal strategies such as FakePolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader].
-Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [zhou2024stealthdiffusion].
+Gradient sign-based attacks [23, 28, 16], latent space manipulations [29], and black-box evaluations [3, 42] reveal vulnerabilities in existing detectors.
+Other works leverage frequency-domain perturbations [25], natural degradation noise [19], or fingerprint removal strategies such as FakePolisher [22] and TraceEvader [56].
+Unlike Deepfakes, which primarily manipulate real face images, AI-generated images encompass far more diverse synthetic content, amplifying disinformation risks [64].
 In this work, we propose a universal attack against both AIGI and Deepfake detectors, evaluating under both white-box and practical black-box settings.
@@
 Eq.
-1 can be performed with iterative gradient-based methods, such as PGD [pgd] or I-FGSM [I-FGSM]:
+1 can be performed with iterative gradient-based methods, such as PGD [36] or I-FGSM [28]:
 $\tilde{\mathbf{x}}^{i+1}$ $=\tilde{\mathbf{x}}^{i}+\alpha\cdot\operatorname{sign}(\nabla L\left(\tilde{\mathbf{x}}^{i},y,\Theta\right))$
@@
 III-B Frequency-based Analysis and Attacks
-Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [wang2020cnn, zhong2023rich].
-While these subtle clues are invisible in the spatial domain, a series of works [frank2020leveraging] demonstrate that there are obvious differences between real and fake images in the frequency domain.
+Many AIGI detection approaches distinguish between real and fake images via subtle artifacts [54, 63].
+While these subtle clues are invisible in the spatial domain, a series of works [14] demonstrate that there are obvious differences between real and fake images in the frequency domain.
 This inspires us to explore the vulnerable region of AIGI detectors from a frequency perspective.
 To this end, we first implement the discrete cosine transform (DCT) $\mathcal{D}(\cdot)$ to transfer the inputs from the spatial domain to the frequency domain.
-To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [long2022frequency] to visualize the sensitive components of real and fake images across different models:
+To investigate the difference between real images and fake images in the frequency domain, we use the spectrum saliency map [34] to visualize the sensitive components of real and fake images across different models:
 $\mathbf{S}_{\Theta}=\frac{\partial J(\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}),y,\Theta)}{\partial\mathcal{D}(\mathbf{x})},$ (3)
@@
 $\Gamma(\mathbf{x})=\mathcal{D}^{-1}(\mathcal{D}(\mathbf{x}+\xi)\odot\mathcal{M}),$ (4)
-where $\Gamma(\cdot)$ denotes the random spectrum transformation [long2022frequency].
+where $\Gamma(\cdot)$ denotes the random spectrum transformation [34].
 $\odot$ is the Hadamard product, $\xi$ is a random noise drawn from an isotropic Gaussian $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$, and each element of $\mathcal{M}$ is sampled from a Uniform distribution $\mathcal{U}(1-p,1+p)$.
@@
 Although tuning the spectrum transformation in Eq.
-5 can simulate different substitute models with a homogeneous architecture [long2022frequency], it shows limited transferability when applied to heterogeneous architectures, e.g., transferring adversarial examples across ViTs and CNNs.
+5 can simulate different substitute models with a homogeneous architecture [34], it shows limited transferability when applied to heterogeneous architectures, e.g., transferring adversarial examples across ViTs and CNNs.
 This motivates us to consider the frequency-based attack from a Bayesian perspective, i.e., exploring the full posterior distribution of the surrogate model to further narrow this gap between heterogeneous models.
@@
 However, it is not straightforward to attack AIGI detectors in such a Bayesian manner due to several factors.
-First, the Bayesian posterior for DNNs is a high-dimensional distribution due to a very large number of parameters of DNNs [izmailov2021bayesian].
+First, the Bayesian posterior for DNNs is a high-dimensional distribution due to a very large number of parameters of DNNs [24].
 Hence computing and sampling the posterior distribution is an intractable problem.
 Albeit feasible for approximately sampling the posterior via variational inference or Markov Chain Monte Carlo (MCMC), it is computationally slow and expensive in such a high-dimensional space.
-Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [zhu2024genimage, he2021forgerynet].
+Furthermore, to improve the accuracy and generalization of the AIGI detectors, there is a growing inclination to train AIGI detectors on large-scale datasets [66, 18].
 From the perspective of end-users, it is not desirable to re-train a surrogate model on large-scale datasets for attack.
@@
 III-C2 Inference on Bayesian Appended Models
-$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[springenberg_bayesian_2016] to sample appended model $\theta^{\prime}$ in each iteration:
+$\Theta$ is frozen after pre-training. We use Stochastic Gradient Adaptive Hamiltonian Monte Carlo[49] to sample appended model $\theta^{\prime}$ in each iteration:
 $\theta^{\prime}_{t+1}=\theta^{\prime}_{t}-\sigma^{2}\mathbf{C}^{-1/2}_{\theta^{\prime}_{t}}\mathbf{h}_{\theta^{\prime}_{t}}+\mathbf{N}(0,2F\sigma^{3}\mathbf{C}^{-1}_{\theta^{\prime}_{t}}-\sigma^{4}\mathbf{I}),$
 $\mathbf{C}_{\theta^{\prime}_{t}}\leftarrow(1-\tau^{-1})\mathbf{C}_{\theta^{\prime}_{t}}+\tau^{-1}\mathbf{h}_{\theta^{\prime}_{t}}^{2},$ (10)
-where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [springenberg_bayesian_2016].
+where $\sigma$ represents the step size, $F$ denotes the friction coefficient, $\mathbf{h}$ is the stochastic gradient of the system, $\mathbf{N}$ represents a Normal distribution, $\mathbf{I}$ stands for an identity matrix, $\mathbf{C}$ is a pre-conditioner updated through an exponential moving average, and $\tau$ is chosen automatically [49].
 III-D Hybrid Adversarial Attack
-Despite detecting fake fingerprints in the frequency domain, some works also extract fingerprint features in the spatial domain [wang2020cnn].
+Despite detecting fake fingerprints in the frequency domain, some works also extract fingerprint features in the spatial domain [54].
 We hence incorporate the attack gradient from the frequency domain with the spatial gradient to further improve the adversarial transferability across different domains.
@@
 Datasets: We chose three generated image datasets created by a wide range of generative models.
-Synthetic LSUN is a commonly used dataset proposed by CNNSpot [wang2020cnn], containing 360k real images from LSUN and 360k fake images generated by ProGAN [progan].
-GenImage [zhu2024genimage] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
-Following the protocol in [zhu2024genimage], we employ a subset of GenImage, collecting 162k real images from Imagenet [deng2009imagenet] and 162k Stable Diffusion(SD) V1.4 [sd] generated images for training.
+Synthetic LSUN is a commonly used dataset proposed by CNNSpot [54], containing 360k real images from LSUN and 360k fake images generated by ProGAN [26].
+GenImage [66] is a recently proposed large-scale dataset, containing 1331k real images and 1350k fake images generated by eight generative models.
+Following the protocol in [66], we employ a subset of GenImage, collecting 162k real images from Imagenet [4] and 162k Stable Diffusion(SD) V1.4 [47] generated images for training.
 The images generated by the other generators are used for testing in Section IV-C2.
-To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [shamshad2023evading].
-The Deepfake dataset consists of 50k real face images from FFHQ [ffhq] and 50k generated face images generated by StyleGAN2 [ffhq].
+To verify our proposed attack is a universal threat across AIGI and Deepfake detection, we also employ the synthetic FFHQ face dataset proposed by [48].
+The Deepfake dataset consists of 50k real face images from FFHQ [27] and 50k generated face images generated by StyleGAN2 [27].
 After training, we collect only the correctly classified testing samples for attack in evaluation.
@@
 Evaluated Models: We extensively evaluate the transferability of adversarial examples by 17 state-of-the-art AIGI detectors, including heterogeneous model architectures and various detection methods.
-For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [wang2020cnn], MobileNet [howard2017mobilenets], EfficientNet [tan2019efficientnet] and DenseNet [huang2017densely], and ViT-based detectors Vision Transformer(ViT) [vit] and Swin-Transformer(Swin-ViT) [liu2021swin].
-For evaluating on various detection methods, we use frequency-based detectors DCTA [frank2020leveraging], Spec [zhang2019detecting], FreqNet [tan2024frequency], and FreqMask [doloriel2024frequency], gradient-based detectors LGrad [tan2023learning], CLIP-based detectors UnivFD [ojha2023towards], and diffusion-based detectors DNF [zhang2023diffusion], DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade].
-In addition, LNP [liu2022detecting] extracts the noise pattern of images and GramNet [GramNet] learns the global texture representation.
+For evaluating on different model architectures, we choose CNN-based detectors CNNSpot [54], MobileNet [20], EfficientNet [52] and DenseNet [21], and ViT-based detectors Vision Transformer(ViT) [11] and Swin-Transformer(Swin-ViT) [32].
+For evaluating on various detection methods, we use frequency-based detectors DCTA [14], Spec [61], FreqNet [50], and FreqMask [9], gradient-based detectors LGrad [51], CLIP-based detectors UnivFD [43], and diffusion-based detectors DNF [62], DIRE [55] and AEROBLADE [46].
+In addition, LNP [31] extracts the noise pattern of images and GramNet [33] learns the global texture representation.
 We also consider them as victim models.
@@
 Note that AEROBLADE is a training-free method and not trained on the listed datasets.
-Compared Methods: We adopt gradient-based methods, I-FGSM [I-FGSM], PGD [pgd], MI [dong2018boosting] and frequency-based methods S2[long2022frequency], SSAH [luo2022frequency] and ensemble-based methods ENS [dong2018boosting] and SVRE [svre].
-Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [jia2022exploring], Fakepolisher [huang2020fakepolisher] and TraceEvader [wu2024traceevader] as baselines.
+Compared Methods: We adopt gradient-based methods, I-FGSM [28], PGD [36], MI [10] and frequency-based methods S2[34], SSAH [35] and ensemble-based methods ENS [10] and SVRE [58].
+Because attacks against Deepfake detectors are the most similar adversaries to ours, we also consider state-of-the-art Deepfake detection attacks FaceAttacker [25], Fakepolisher [22] and TraceEvader [56] as baselines.
 Fakepolisher and TraceEvader remove the Deepfake traces and we follow their default settings.
 For other iterative attacks, we run 10 iterations with step size $\alpha=2/255$ under $l_{\infty}$ perturbation budget of $8/255$ for all these attacks.
-Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [wang2020cnn].
+Implementation Details: For MobileNet, EfficientNet, DenseNet, ViT and Swin-Transformer(Swin-ViT), we train them following the default setting in CNNSpot [54].
 Specifically, we choose classifiers pre-trained on ImageNet, and train them with Adam optimizer using Binary Cross-Entropy loss with an initial learning rate of 0.0001.
-For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [wang2020cnn] to improve the models’ generalization and robustness.
+For a fair evaluation, we follow the same data augmentation strategy used in CNNSpot [54] to improve the models’ generalization and robustness.
 Before cropping, images are blurred with $\sigma\sim\text{Uniform}[0,3]$ with 10% probability, and JPEG-ed with 10% probability.
-We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [wang2020cnn].
+We transformed images to 224 pixels on the Synthetic LSUN(ProGAN) and GenImage(SD) datasets following CNNSpot [54].
 On the FFHQ (StyleGAN2) dataset, we resized images to 224 pixels to ensure the integrity of the real/fake faces.
@@
 For other AIGI detectors, we use the pre-trained model from their official code.
-For post-train Bayesian optimization, we follow the default setting in [bbc].
+For post-train Bayesian optimization, we follow the default setting in [53].
 Although BNNs theoretically necessitate sampling numerous for inference, in practice, we find the number of models $K=3$ is adequate.
 Opting for a larger number of appended models escalates computational overhead; thus, we opt for $K=3$.
-For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [long2022frequency].
+For frequency-based attack, we set the tuning factor $\rho$ = 0.5 for $\mathcal{M}$, the standard deviation $\sigma$ of $\xi$ is set to the value of $\epsilon$, following [34].
 All experiments were conducted on 4 NVIDIA GeForce RTX 3090s.
@@
 FPBA4 98.9 98.0 76.4 95.9 19.8 48.0 51.5 94.8 69.2
-TABLE V: Transfer-based attack against SOTA detectors on Synthetic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [wang2020cnn].
+TABLE V: Transfer-based attack against SOTA detectors on Synthetic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [54].
 Attack GramNet LGrad LNP UnivFD DNF FreqNet FreqMask Ave
@@
 These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors.
-For example, techniques like searching for adversarial points on the face manifold [li2021exploring], removing facial forgery traces [huang2020fakepolisher, wu2024traceevader] or simply updating perturbations in the frequency domain to improve attack imperceptibility [jia2022exploring] may not be applicable for AIGI detectors.
-To demonstrate this, we compare with FakePolisher [huang2020fakepolisher], TraceEvader [wu2024traceevader] and FaceAttacker [jia2022exploring], which are the SOTA attacks for Deepfake detection.
+For example, techniques like searching for adversarial points on the face manifold [29], removing facial forgery traces [22, 56] or simply updating perturbations in the frequency domain to improve attack imperceptibility [25] may not be applicable for AIGI detectors.
+To demonstrate this, we compare with FakePolisher [22], TraceEvader [56] and FaceAttacker [25], which are the SOTA attacks for Deepfake detection.
 Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images.
@@
 IV-B2 Comparison with Ensemble-based Attacks
-Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [dong2018boosting] and SVRE [svre], i.e., utilizing an ensemble of surrogate models to generate adversarial examples.
+Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [10] and SVRE [58], i.e., utilizing an ensemble of surrogate models to generate adversarial examples.
 To examine the impact of different surrogate combinations(CNN-based vs.
@@
 Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts.
-Specifically, GramNet [GramNet] extracts the global texture representation; UnivFD [ojha2023towards] captures semantic feature using pre-trained CLIP [CLIP]; LNP [liu2022detecting] extracts the noise pattern of images; LGrad [tan2023learning] extract gradient information through a pre-trained model; DNF [zhang2023diffusion] leverages diffusion noise features, extracting them through an inverse diffusion process [adm]; FreqNet [tan2024frequency] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [doloriel2024frequency] introduces a frequency-domain mask strategy for data augmentation during training.
+Specifically, GramNet [33] extracts the global texture representation; UnivFD [43] captures semantic feature using pre-trained CLIP [44]; LNP [31] extracts the noise pattern of images; LGrad [51] extract gradient information through a pre-trained model; DNF [62] leverages diffusion noise features, extracting them through an inverse diffusion process [5]; FreqNet [50] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [9] introduces a frequency-domain mask strategy for data augmentation during training.
 These significant differences among various detection methods present challenges for the transferability of adversarial examples.
@@
 FPBA 43.1 40.8
-Very recently, reconstruction error-based methods, such as DIRE [wang2023dire] and AEROBLADE [ricker2024aeroblade] have emerged as competitive approaches for detecting diffusion-generated images.
+Very recently, reconstruction error-based methods, such as DIRE [55] and AEROBLADE [46] have emerged as competitive approaches for detecting diffusion-generated images.
 Therefore, we adopt them as victim models in our evaluation.
@@
 Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX.
-We follow [ricker2024aeroblade] to report detection performance using average precision (AP).
+We follow [46] to report detection performance using average precision (AP).
 FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
@@
 We hence evaluate the robustness of cross-generator image detection to investigate whether adversarial examples are a real threat to AIGI detection.
-We train the detectors on images generated by SD v1.4 [sd] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [Midjourney], SD V1.4 [sd], SD V1.5 [sd], ADM [adm], Wukong [wukong] and BigGAN [biggan].
-Because Swin-ViT achieves the SOTA results on different subsets [zhu2024genimage], we use it as the surrogate model.
+We train the detectors on images generated by SD v1.4 [47] and assess their robustness against adversarial examples, in which the adversarial perturbations are added to the images generated by Midjourney [41], SD V1.4 [47], SD V1.5 [47], ADM [5], Wukong [57] and BigGAN [2].
+Because Swin-ViT achieves the SOTA results on different subsets [66], we use it as the surrogate model.
 The benign accuracy and attack performance on unseen source data are reported in Table VI and Table VII respectively.
@@
 We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples.
-We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [yanorthogonal].
+We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [59].
 As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack.
@@
 As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors.
-In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [pgd] and TRADES [trades].
+In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [36] and TRADES [60].
 However, we found that the training process of AT fails to converge.
@@
 Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies.
-Considering that AIGI detection typically [wang2020cnn] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [wang2020cnn, zhu2024genimage] to utilize Gaussian blurring and JPEG compression as a defense.
+Considering that AIGI detection typically [54] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [54, 66] to utilize Gaussian blurring and JPEG compression as a defense.
 In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios.
@@
 However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under real-world scenarios.
-Next, we follow the suggestions from [wang2020cnn, zhu2024genimage] to simultaneously utilize Gaussian blurring and JPEG compression as a defense.
+Next, we follow the suggestions from [54, 66] to simultaneously utilize Gaussian blurring and JPEG compression as a defense.
 As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@
 IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
-In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have a higher success rate [dong2018boosting].
-It indicates the possibility of gradient masking [athalye2018obfuscated].
+In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have a higher success rate [10].
+It indicates the possibility of gradient masking [1].
 To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties.
@@
 But we empirically found that FPBA is also effective for face forgery detection.
-To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [shamshad2023evading], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [ffhq].
+To demonstrate this, we conduct the experiments on synthetic FFHQ face datasets [48], which consists of 50k real face images from FFHQ and 50k fake generated face images created by StyleGAN2 [27].
 We report the results in Table XIII.
@@
 This work was supported in part by the NSF China (No. 62302139, 62576020) and Fundamental Research Funds for the Central Universities of China (PA2025IISL0113, JZ2025HGTB0227).
-[Uncaptioned image] Yunfeng Diao is a Lecturer in the School of Computer Science and Information Engineering, Hefei University of Technology, China.
-He received his PhD from Southwest Jiaotong University, China.
-His current research interests include computer vision and the security of machine learning.
-He has published over 30 papers in leading venues such as CVPR, ICLR, ICCV, ICML, IEEE TMM and IEEE TCSVT.
-He has been an editorial board member of IJAACS, the chief organizer of several workshops at IJCAI, also regularly reviewing for top-tier journals and conferences.
-He has received the Outstanding Contribution Award at the IJCAI Workshop and the Best Paper Award at the IROS RODGE Workshop.
-[Uncaptioned image] Naixin Zhai received the B.S.
-degree from Hefei University of Technology, Hefei, China, in 2024.
-He is current working toward the M.S.
-degree with the Department of Electronic Engineering and Information Science.
-His research interests include deep learning and synthetic image detection.
-[Uncaptioned image] Changtao Miao received his B.S.
-degree in 2019 from AnHui University.
-He is currently pursuing the Ph.D.
-degree in Cyber Science and Technology in University of Science and Technology of China.
-His research interests include face forgery forensics and face manipulation.
-[Uncaptioned image] Zitong Yu (Senior Member, IEEE) received the Ph.D.
-degree in computer science and engineering from the University of Oulu, Finland, in 2022.
-Currently, he is an Assistant Professor with Great Bay University, China.
-He was a Post-Doctoral Researcher with the ROSE Laboratory, Nanyang Technological University.
-He was a Visiting Scholar with TVG, University of Oxford, from July 2021 to November 2021.
-His research interests include human-centric computer vision and biometric security.
-He was a recipient of IAPR Best Student Paper Award, IEEE Finland Section Best Student Conference Paper Award 2020.
-[Uncaptioned image] Xingxing Wei received his Ph.D degree in computer science from Tianjin University, and B.S.
-degree in Automation from Beihang University (BUAA), China.
-He is now an Associate Professor at Beihang University (BUAA).
-His research interests include computer vision, adversarial machine learning and its applications to multimedia content analysis.
-He is the author of referred journals and conferences in IEEE TPAMI, TMM, TCYB, TGRS, IJCV, PR, CVIU, CVPR, ICCV, ECCV, ACMMM, AAAI, IJCAI etc.
-[Uncaptioned image] Xun Yang (Senior Member, IEEE) received the Ph.D.
-degree from the Hefei University of Technology, Hefei, China, in 2017.
-He is currently a Professor with the Department of Electronic Engineering and Information Science, University of Science and Technology of China (USTC).
-From 2015 to 2017, he visited the University of Technology Sydney (UTS), Australia, as a Joint Ph.D.
-Student.
-He was a Research Fellow with the NExT++ Research Center, National University of Singapore (NUS), from 2018 to 2021.
-His current research interests include information retrieval, cross-media analysis and reasoning, and computer vision.
-He serves as an Associate Editor for IEEE TRANSACTIONS ON BIG DATA and Multimedia Systems journal.
-[Uncaptioned image] Meng Wang (Fellow, IEEE) is a professor at Hefei University of Technology, China.
-His current research interests include multimedia content analysis, computer vision, and pattern recognition.
-He received paper prizes or awards from ACM MM 2009 (Best Paper Award), ACM MM 2010 (Best Paper Award), MMM 2010 (Best Paper Award), SIGIR 2015 (Best Paper Honorable Mention), IEEE TMM 2015 and 2016 (Prize Paper Award Honorable Mention) and ACM TOMM 2018 (Nicolas D.
-Georganas Best Paper Award), etc.
-He currently serves on the editorial/advisory boards of IEEE TPAMI, IEEE TMM, IEEE TNNLS, etc.
-He is a fellow of IEEE and IAPR.
-Vulnerabilities in AI-generated Image Detection:
-The Challenge of Adversarial Attacks
-—Supplementary Document—
-Figure 10: Visualization of the spectrum saliency map with different models.
-VI Additional Experiments
-VI-A Experiment Details
-We compared a series of distinguished backbone architectures pretrained on the ImageNet dataset.
-Specifically, we evaluated ResNet-50, DenseNet-121, EfficientNet-B4, MobileNet-V2, ViT-B/16, and Swin-B.
-Additionally, we examined Spec equipped with ResNet-34 architecture, as well as DCTA equipped with ResNet-50 architecture.
-In Table XVIII, we enumerate the accuracy of all models on the test set as observed in our experiments.
-The meticulous evaluation of these foundational models not only revealed the potential vulnerabilities of AIGI detection models but also highlighted the adversarial threats they face.
-Our analytical results offer significant insights for understanding and improving the robustness of AIGI detection models.
-TABLE XVIII: The accuracy of each detector on different datasets.
-Model ResNet DenNet EffNet MobNet Spec DCTA ViT Swin
-LSUN(ProGAN) 99.99 99.91 99.96 99.93 99.90 99.91 99.74 100.00
-GenImage(SD) 99.55 99.92 99.9 99.93 99.12 99.17 99.48 99.97
-FFHQ(Style GAN) 99.81 99.89 99.83 99.63 96.51 96.6 99.55 99.99
-VI-B Evaluation on ResNeXt
-We use ResNext [xie2017aggregated] as the surrogate to evaluate the attack performance.
-As reported in Table XIX, FPBA still achieves the best average attack success rate, which is consistent with other surrogate models.
-TABLE XIX: Attack evaluation using ResNeXt as surrogate model on the ProGAN-LSUN dataset.
-Attack ResNeXt ResNet DenNet Spec DCTA ViT Swin Ave.
-MI-FGSM 54.2 29.7 21.3 24.8 49.7 14.2 6.6 28.6
-PGD 77.1 35.4 30.9 28.0 50.4 20.6 28.0 38.6
-S2I 98.6 90.6 86.9 13.4 36.9 15.0 68.0 58.5
-FPBA 99.9 89.0 89.2 21.5 49.0 29.6 64.1 63.2
-VI-C Study of Attack Strength
-As shown in Table XX, we report the comparison of the attack success rates of our FPBA with existing methods under different attack strengths.
-The results reflect a pattern consistent with when $\epsilon$=8/255, where our method can achieve globally optimal results compared to previous methods, and a larger perturbation budget consistently enhances adversarial transferability across various surrogate models.
-However, as the attack strength increases, the noise added causes more and more damage to the image, and the visual impact becomes greater, which is revealed in Fig.
-11.
-VI-D Spectrum Saliency Map
-In Fig. 10, we present the spectrum saliency map of all detectors for the original samples and the spectrum saliency map after frequency domain transformation.
-TABLE $X$: The attack success rate(%) on CNN-based, Vit-based and Frequency-based models on the Synthetic LSUN(ProGAN) with $\epsilon$=4/255, 8/255 or 12/255.
-“Average” was calculated as the average transfer success rate over all victim models except for the surrogate model.
-Surrogate Model Attack Methods ResNet DenNet EffNet MobNet Spec DCTA ViT Swin Average
-IFGSM 52.1 44.6 14.0 36.8 15.0 41.2 12.9 24.6 27.0
-MIFGSM 52.1 50.0 23.5 45.4 15.0 45.1 22.9 38.3 34.3
-ResNet-50 PGD 53.5 46.3 15.5 38.7 20.3 44.4 12.4 26.0 29.1
-S2I 97.1 68.1 21.1 33.0 7.0 18.3 2.0 17.9 23.0
-$\epsilon$=4/255 FPBA(Ours) 98.9 93.7 55.0 79.8 15.8 40.8 18.6 75.1 54.1
-IFGSM 52.1 51.3 24.6 48.4 17.7 48.0 35.2 46.2 38.7
-MIFGSM 52.1 51.5 27.9 47.6 26.1 49.8 37.5 42.7 40.4
-ResNet-50 PGD 78.3 77.6 49.4 73 25.6 47.6 41.0 70.2 54.9
-S2I 97.8 86.4 61.5 78.6 20.5 40.8 11.7 74.5 53.4
-$\epsilon$=8/255 FPBA(Ours) 98.9 98.0 76.4 95.9 19.8 48 51.5 94.8 69.2
-IFGSM 52.1 51.9 28.7 50.0 18.3 49.1 40.6 50.2 41.2
-MIFGSM 52.1 51.8 28.7 47.4 37.2 50.0 34.7 38.9 41.2
-ResNet-50 PGD 96.3 96.1 67.0 91.9 30.6 48.8 58.6 87.5 68.6
-S2I 97.3 88.4 78.5 89.3 28.5 47.6 25.4 85.3 63.3
-$\epsilon$=12/255 FPBA(Ours) 98.9 99.3 82.9 98.5 20.7 49.2 68.7 98.0 73.9
-IFGSM 4.0 5.6 12.5 75.4 12.3 24.6 4.8 3.4 9.6
-MIFGSM 9.0 11.5 16.5 75.4 13.7 26.4 6.4 10.9 13.5
-MobileNet-V2 PGD 3.4 5.0 11.8 87.7 15.1 25.7 4.5 3.2 9.8
-S2I 6.0 7.0 19.2 93.6 5.9 8.0 1.2 5.6 7.6
-$\epsilon$=4/255 FPBA(Ours) 7.9 11.8 26.0 99.5 14.4 25.8 4.5 9.5 14.3
-IFGSM 14.3 17.8 22.4 75.4 17.2 34.9 9.2 18.4 19.2
-MIFGSM 23.8 23.9 28.6 75.4 19.6 42 15.5 24.8 25.5
-MobileNet-V2 PGD 20.3 26.0 20.6 97.6 25.0 40.4 9.7 29.9 24.6
-S2I 14.4 17.5 34.7 97.8 16.3 25.4 4.3 14.6 18.2
-$\epsilon$=8/255 FPBA(Ours) 32.2 40.4 51.7 99.6 22.6 36.4 12.3 44.6 34.3
-IFGSM 18.8 21.4 27.7 75.4 18.0 40.2 11.4 23.0 22.9
-MIFGSM 25.2 25.2 33.7 75.4 22.5 48.7 21.6 25.2 28.9
-MobileNet-V2 PGD 34.8 39.8 27.6 99.8 35.4 48.5 14.7 45.1 35.1
-S2I 28.1 33.5 44.0 98.6 23.3 34.5 7.9 35.5 29.5
-$\epsilon$=12/255 FPBA(Ours) 42.3 45.7 56.3 99.9 27.0 42.0 18.7 49.0 40.1
-Figure 11: The visual results of adversarial examples with different perturbation budgets.